### v1

# MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

Abstract
*Equal Contribution.
Corresponding author.
Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample. In this paper, we introduce score-based iterative reconstruction (SIR), an efficient and general algorithm for 3D generation with a multi-view score-based diffusion model. Given the images produced by the diffusion model, SIR reduces NFEs by repeatedly optimizing 3D parameters, unlike the single optimization in SDS, mimicking the 3D reconstruction process. With other improvements including optimization in the pixel space, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks. In particular, retaining a comparable performance, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field and takes about 20 seconds to generate meshes from 3D Gaussian splitting on a single A100 GPU, halving the time of the fastest zero-shot baseline, DreamGaussian. Our code is available at https://github.com/ML-GSAI/MicroDreamer.
Keywords:
3D Generation Diffusion Model Multi-view Diffusion Score Distillation Sampling Zero-shot
Figure 1: MicroDreamer can generate high-quality meshes in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data.
1 Introduction
Recently, optimization-based approaches [40, 56, 19, 66, 5, 24, 46, 60, 55, 18, 53, 33, 62] particularly score distillation sampling (SDS) [40, 56] have emerged as promising avenues for 3D generation based on text-to-image diffusion models [48, 11, 51, 43, 44, 42]. These approaches are appealing due to their minimal, or even zero reliance on 3D data, in contrast to the data-intensive requirements of feed-forward approaches [3, 4, 6, 14, 37, 26, 13, 35, 59, 61, 65, 54, 23, 22, 67]. This advantage is particularly significant given that 3D data are costly and scarce.
Despite their promising capabilities, optimization-based approaches suffer from low efficiency. Intrinsically, the inefficiency lies in the extensive number of function evaluations (NFEs), i.e. forward passes of the diffusion model, required for each 3D object generation. Besides, within the commonly employed latent diffusion model (LDM) [43], the gradients of the 3D parameters need to be backpropagated through an encoder that maps rendered images to the latent space, further exacerbating the inefficiency of these approaches.
To address the computational issues, this paper introduces an efficient and general algorithm termed score-based iterative reconstruction (SIR) for zero-shot 3D generation in Sec. 3. Similar to SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data. However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a single set of images produced by the diffusion model, mimicking the 3D reconstruction process to reduce the total NFEs (see Fig. 3(a)). To obtain 3D-consistent and high-quality images as the ground truth for 3D reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration. Besides, even mapped back to the pixel space through the decoder in LDM, the refined images are still of high quality for reconstruction, enabling optimization in the pixel space to further speed up (see Fig. 6(a)-6(b)).
We delve into critical components beyond the algorithm to ensure compatibility with SIR. The comprehensive system is named by MicroDreamer, highlighting its efficiency for 3D generation. As detailed in Sec. 4, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3D Gaussian splatting (3DGS)[15].
SIR and MicroDreamer are broadly applicable to both neural radiance fields (NeRF)[34] and 3DGS, and to both text-to-3D and image-to-3D tasks, as detailed in Sec. 5. Employing three widely adopted multi-view diffusion models [47, 52, 58], we systematically compare SIR and SDS for NeRF generation. Retaining a comparable performance, SIR can accelerate the generation process by 5 to 20 times. Besides, MicroDreamer can efficiently produce 3DGS and further refine it into high-quality meshes, delivering consistent 3D meshes in about 20 seconds on a single A100 GPU – about twice as fast as the most competitive optimization-based alternatives, DreamGaussian [55]. Remarkably, MicroDreamer is on par with speed compared to feed-forward methods trained on extensive 3D data, with a very competitive CLIP similarity [41, 55].
Figure 2: Overview of SIR and SDS. $\mathcal{E}$, $\mathcal{D}$ and $\epsilon_{\phi}$ are the encoder, decoder, and diffusion model in LDM [43]. The solid and dashed arrows represent forward and backward passes respectively. The computation bottlenecks are highlighted in red. SIR is efficient because it lowers the NFEs in total and enables optimization in the pixel space.
2 Background
We present background on 3D representations, diffusion models, multi-view diffusion models, and current optimization-based 3D generation methods.
2.1 3D Representation
Neural radiance fields[34, 36] (NeRF) and 3D Gaussian splatting[15] (3DGS) have emerged as popular 3D representations. NeRF employs an MLP to predict the colors and density of the input space coordinates. 3DGS consists of multiple 3D Gaussians parameterized by the colors, center, scale, and rotation quaternion. We denote the corresponding tunable parameters in both representations as $\theta$. Given camera poses $c$, both approaches define a differential rendering process, denoted by $g(\theta,c)$. They are proven efficient and effective in 3D reconstruction [34, 15] and generation [40, 55].
2.2 Diffusion Models
A diffusion model [48, 11, 51] consists of a forward process and a sampling process. The forward process gradually adds Gaussian noise to an input image from time 00 to $T$. For any $t\in(0,T)$, the noise-adding process can be written as follows:
$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon:=\textrm{Noise-adding}(x_{0},0\rightarrow t),\quad\epsilon\in\mathcal{N}(0,I),$ (1)
where the coefficients $\alpha_{t}$ and $\sigma_{t}$ form the noise schedule of the diffusion model. A noise prediction network $\epsilon_{\phi}(x_{t},t)$ with parameters $\phi$ is trained to predict the noise in the input $x_{t}$ with the corresponding noise level, i.e. time $t$. Plugging in the noise prediction network into Eq. (1), we can solve $x_{0}$ from a noisy image $x_{t}$ of time $t$ by a single-step prediction as follows:
$\hat{x}^{t}_{0}=\frac{1}{{\alpha}_{t}}x_{t}-{\frac{\sigma_{t}}{{\alpha}_{t}}}\epsilon_{\phi}(x_{t},t),$ (2)
which is an efficient approximation of the original input.
Instead of using Eq. (2) directly, the sampling process of diffusion gradually denoises through the noise prediction network and generates images from pure Gaussian noise. Among existing samplers [21, 2, 1, 30, 31, 64], the denoising diffusion implicit models (DDIM) [49] facilitate a sequence of samplers with random noise control $\eta$. When $\eta=0$, it solves the equivalent probability flow ordinary differential equation (ODE) [51] of the diffusion model and enjoys a fast sampling process with fewer number of function evaluations (NFEs) [^1] Throughout the paper, we refer to the number of forward passes through $\epsilon_{\phi}$ as NFEs.. In this setting, the one-step sampling of DDIM is given by:
$x_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}+\left(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Sampler}}(x_{t},t\rightarrow t-1).$ (3)
We refer the readers to the original paper [49] for the formula of $\eta>0$. For $\eta=1$, it represents a standard SDE sampler [11], with a higher tolerance for mismatches in the distributions of latent variables [29, 39].
Besides, when $\eta=0$ there is an inverse process (called DDIM inversion) that maps the distribution of images to the same distributions of noisy images as in Eq. (1) but it can maintain the unique feature of an input image and reconstruct it accurately through the corresponding DDIM sampler in Eq. (3). The (one-step) DDIM inversion [49] is formulated as follows:
$x_{t+1}=\frac{\alpha_{t+1}}{\alpha_{t}}x_{t}+\left(\sigma_{t+1}-\frac{\alpha_{t+1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Inversion}}(x_{t},t\rightarrow t+1).$ (4)
2.3 Multi-view Diffusion Models
After being trained on a modest amount of 3D data, diffusion models can generate 3D-consistent multi-view, known as multi-view diffusion models. Among them, MVDream [47] takes text inputs and outputs multi-view images consistent in 3D. In contrast, Zero-1-to-3 [24] and ImageDream [58] focus on the Image-to-3D task, taking an additional reference image and its viewpoint as input. These models output new viewpoint images consistent with the reference image. This paper directly utilizes these pre-trained multi-view diffusion models without any further fine-tuning. For more details, please refer to the original papers and the source code in our supplementary materials. Notably, such multi-view diffusion models cannot provide a sufficient number of consistent multi-view images for classical 3D reconstruction, as shown in Fig. 6(c).
2.4 Optimization-based Algorithms for 3D Generation
Built upon (multi-view) diffusion models, optimization-based methods (see Sec. 6 for a review) aim to generate 3D content in a zero-shot manner. Among them, score distillation sampling (SDS) [40, 56] is the most representative and popular approach. Formally, denoting the rendered images as $x=g(\theta,c)$, SDS repeats adding noise to $x$ according to Eq. (1) and updating the 3D parameters $\theta$ by
$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi):=\mathbb{E}_{t}\left[w(t)(\epsilon_{\phi}(\alpha_{t}x+\sigma_{t}\epsilon,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (5)
where $w(t)$ is a fixed weighting function and we omit the dependency on the prompt $y$ and camera $c$ for simplicity. Notably, by reparameterization [66] according to Eq. (2), SDS has an equivalent data-prediction form:
$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi)=\mathbb{E}_{t}\left[\frac{w(t)\alpha_{t}}{\sigma_{t}}(x-\hat{x}_{0}^{t})\frac{\partial x}{\partial\theta}\right].$ (6)
3 Score-based Iterative Reconstruction
We present the general framework of score-based iterative reconstruction (SIR), an efficient and versatile algorithm inspired by classical 3D reconstruction in Sec. 3.1. SIR is designed to minimize the number of function evaluations (NFEs) typically required by existing methods. In Sec. 3.2, we introduce a diffusion-based process to produce multi-view images as ground truth for reconstruction. In Sec. 3.3, we demonstrate how SIR facilitates the optimization of 3D content directly within pixel space, thereby further boosting efficiency.
3.1 Mimicking 3D Reconstruction to Reduce the NFEs
Existing zero-shot 3D generation algorithms [40, 56, 66, 60] iteratively optimize the 3D parameters based on a 2D diffusion (see Eq. (5) for SDS as an example). Consequently, these methods necessitate a high number of function evaluations (NFEs), presenting a critical challenge to computational efficiency.
Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of traditional 3D reconstruction—could substantially decrease the overall NFEs required. However, traditional 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [24, 47, 58, 52] (see Sec. 6 for a review and discussion). Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple rounds of refinement, dubbed score-based iterative reconstruction (SIR), detailed as follows.
Formally, SIR consists of $K$ reconstruction iterations. In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given an initial 3D parameter $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec. 4) and employ a rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:
$x^{(k)}=g(\theta^{(k)}_{0},c^{(k)}).$ (7)
For simplicity, let $x^{(k)}$ be a vector by flattening all images and concatenating them together and so do any subsequent set of multi-view images.
As detailed in Sec. 3.2, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [47, 24, 58, 52]. The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for the current iteration. In particular, starting from $\theta_{0}^{(k)}$, we optimize the 3D parameters w.r.t. the following reconstruction loss for $I^{(k)}$ steps:
$\mathcal{J}_{\textrm{SIR}}(\theta;c^{(k)},\hat{x}^{(k)})=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|,$ (8)
[…]
Here we assume the rendered images and those sampled by the diffusion model are the same dimensional for simplicity. We will discuss how to deal with the latent diffusion model (LDM) [43] in Sec. 3.3. Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ in Sec. 3.2.
3.2 Obtaining Multi-view Images for Reconstruction by Diffusion
We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process.
The diffusion model has two equivalent forward processes [51]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq. (1) and the inversion process based on the probability flow ordinary differential equations (ODEs) in Eq. (4). In comparison, the noise-adding process is more efficient without function evaluation, and it aids in eliminating low-quality details through high-frequency noise. Conversely, the inversion process can preserve the 3D consistency and overall information of the current 3D object. Observations from existing experiments [38] indicate that diffusion models tend to generate details closer to time 00 and overall structure as they approach time $T$. Inspired by this, we design a hybrid forward process that initially adds noise to time $t_{1}^{(k)}$ and then perform DDIM inversion to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise.. In contrast to random sampled $t_{2}^{(k)}$ in existing algorithms including SDS [40], we adopt a linearly decreased schedule for $t_{2}^{(k)}$ as $k$ progresses, detailed in Sec. 4. Formally, the process is defined as:
$\tilde{x}^{(k)}=\textrm{Inversion}(\textrm{Noise-adding}(x^{(k)},0\rightarrow t_{1}^{(k)}),t_{1}^{(k)}\rightarrow t_{2}^{(k)}),$ (9)
[…]
Note that generally $x^{(k)}$ does not follow the model distribution defined by the diffusion and the resulting $\tilde{x}^{(k)}$ do not strictly adhere to the marginal distribution at the corresponding time. Nevertheless, we still use existing sampling algorithms to obtain refined images $\hat{x}^{(k)}$ from $\tilde{x}^{(k)}$ as follows:
$\hat{x}^{(k)}={\textrm{Sampler}}(\tilde{x}^{(k)},t_{2}^{(k)}\rightarrow 0).$ (10)
Although other advanced sampling methods [21, 2, 1, 30, 31, 64] exist, we choose the popular DDIM in Eq. (3) and tune its noise hyperparameter $\eta$. The optimal value of $\eta$ depends on the base model and we search for the best one in $\{0,0.5,1\}$ as detailed in Tab. 1. The whole process is presented in Algorithm 1.
Compared with existing methods such as SDS, SIR leverages several NFEs for multiple updates of 3D parameters and lowers the NFEs in total (see Fig. 3(a)).
1
2Input: The number of iterations $K$, an initial 3D object $\theta_{0}^{(0)}$, a set of reconstruction steps $\{I^{(k)}\}_{k=0}^{K-1}$ and a set of camera poses $\{c^{(k)}\}_{k=0}^{K-1}$.
3 Output: A final 3D content $\theta^{(K)}_{0}$. for $k$ from 00 to $K-1$ do
4       Render $N$ images $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)})$ ;
5       Obtain noisy $\tilde{x}^{(k)}$ from forward process for $x^{(k)}$ using Eq. (9);
6       Obtain refined $\hat{x}^{(k)}$ from sampling process for $\tilde{x}^{(k)}$ using Eq. (10);
7       for $i$ from 00 to $I^{(k)}-1$ do
8             Compute loss $L=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|$ using Eq. (8);
9             Compute the gradient $\nabla L$, update $\theta^{(k)}_{i}$ to $\theta^{(k)}_{i+1}$;
10
11       end for
12      Obtain $\theta^{(k+1)}_{0}=\theta^{(k)}_{I^{(k)}}$.
13 end for
Algorithm 1 Score-based Iterative Reconstruction
3.3 Enabling Optimization in the Pixel Space
Within the widely adopted framework of LDM [43], the rendered images are mapped through an encoder $\mathcal{E}$ [16] to a latent space, where the loss function is calculated. Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation. In such a case, the SIR loss in the latent space is given by:
$\mathcal{J}_{\textrm{SIR-latent}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|\mathcal{E}(g(\theta,c^{(k)}))-\hat{x}^{(k)}\|.$ (11)
An alternative approach maps the diffusion output back to the pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space. This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
However, we argue that this is not feasible for existing optimization methods [40, 56, 60] such as SDS. Even reparameterized as the data predicting form in Eq. (6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix A). In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space. The SIR loss in the pixel space is formalized as:
$\mathcal{J}_{\textrm{SIR-pixel}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|g(\theta,c^{(k)})-\mathcal{D}(\hat{x}^{(k)})\|,$ (12)
which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as shown in Fig. 6(a)-6(b). It is set as the default loss throughout the paper.
4 MicroDreamer
In addition to our algorithm SIR, we carefully analyze and design other critical elements in the diffusion model and 3D training. Our comprehensive system is called MicroDreamer to highlight its efficiency.
3D initialization and camera views.
We utilize a direct reconstruction approach to initialize the 3D assets. Specifically, we optimize the loss function in Eq. (8) by several steps (see Tab. 1 for detailed values) to update the 3D content, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion model. The camera views are uniformly sampled following the corresponding baselines [55, 9] except that the azimuth angles of different views in the same iteration are evenly distributed.
Annealed time schedule.
We utilize an annealed time schedule $\{t_{2}^{(k)}\}_{k=0}^{K-1}$ for the end of the forward process. Intuitively, as the quality of the 3D assets improves, the input $x$ in Eq. (7) becomes more proximate to the model distribution, thus requiring fewer steps to refine. This differs from SDS, which samples uniformly random $t_{2}$. Our preliminary experiments in Fig. 6(e) suggest that a linearly annealed schedule is sufficient. The endpoints of the schedule depend on the diffusion model, detailed in Tab. 1. See Appendix B for the schedule of $t_{1}$.
Reference image loss for image-to-3D.
A reference image is available in image-to-3D, which is regarded as the ground truth front view of the 3D object on Stable Zero123[24, 52] following DreamGaussian [55]. In this way, we add the reference loss in the same form as the reconstruction loss with Eq. (8) in each iteration of training. The weight set for the reference loss can be seen in Tab. 1.
3DGS settings and mesh refinement.
For simplicity, our approach to 3DGS largely adheres to the settings in DreamGaussian[55] unless specified. We incorporate a densify and prune procedure at every 100 updates during the initial 300 updates. We follow the mesh extraction method in DreamGaussian [55] and employ a threshold value of $0.2$ in the Marching Cube algorithm [28, 55]. We utilize SIR to optimize the mesh texture in 3 iterations, with 15 steps per iteration. We use a noise-adding process and DDIM with $\eta=0$ for simplicity.
Table 1: Key hyperparameters of MicroDreamer on three base diffusion models. All models are employed to generate NeRF and the last two are employed to generate 3DGS and mesh. The hyperparameters are shared across the two settings by default. Otherwise, those for 3DGS are shown in brackets. $T$ is end time of diffusion.
Model Select MVDream [47] Stable Zero123 [24, 52] ImageDream [58]
Diffusion
CFG [12] 7.5 3.0 3.0 (2.0)
Forward process hybrid hybrid hybrid (random noise only)
Time schedule of $t_{2}$ $0.8T\rightarrow 0.5T$ $0.8T\rightarrow 0.2T$ $0.8T\rightarrow 0.6T$ ($0.8T\rightarrow 0.4T$)
Sampling process DDIM, $\eta=0.0$ DDIM, $\eta=0.5$ DDIM, $\eta=1.0$
Discretization steps 50 20 10
3D training
Resolution $64\rightarrow 128$ $64\rightarrow 128\rightarrow 196$ (256) $64\rightarrow 128$ (256)
Background learned by NN always white always white
# camera views 4 4 (6) 4
# initialized steps 50 15 50
# iterations $K$ 50 30 (20 or 30) 30
# reconstuction steps $I^{(k)}$ 15 15 15
Loss type $\ell_{1}$ $\ell_{1}$ $\ell_{1}$
Reference color loss - 0.1 (0.3) 0
Reference opacity loss - 0.001 (0.01) 0
5 Experiment
We present the results on NeRF, 3DGS, and ablation sequentially. For a fair comparison, we utilize the widely used CLIP similarity [41] for quantitative comparison following DreamGaussian [55]. We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0. For all methods, the corresponding images are rendered from NeRF in Sec. 5.1 and mesh in Sec. 5.2. We report the average CLIP similarities between these images and the reference image or text. See Appendix C for details of the checkpoint and codebase.
5.1 Results on NeRF
We apply SIR algorithm on NeRF[34, 36, 9] leveraging three multi-view diffusion model from MVDream[47], Stable Zero123[52, 24], and ImageDream[58]. For each model, we selected 6 input prompts from a widely used codebase [9] for testing, calculated NFEs, and recorded the change of the average CLIP similarity in the generation process. Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig. 3. Besides, SIR can generate NeRF of high visual quality, detailed in Appendix D, and the conclusion agrees with the results in Fig. 3.
(a) Comparsion w.r.t. NFEs.
(b) Comparison on MVDream[47]
(c) Comparison on Stable Zero123[52, 24]
(d) Comparison on ImageDream[58]
Figure 3: Comparison of SIR and SDS on NeRF. (a) shows that SIR lowers the total NFEs. (b-d) show the change of the CLIP similarity in the generation process under different models and SIR achieves a 5-20 times acceleration approximately.
5.2 Results on 3D Gaussian Splatting
Qualitative comparisons.
We show the generated 3D objects amongst various methods and ours in Fig. 4. The baselines include representative feed-forward models One-2-3-45 [23] and TriplaneGaussian [67], and the fastest optimization-based model DreamGaussian [55]. We present the exported mesh from these models. MicroDreamer generates higher quality mesh exported from 3DGS than the feed-forward baselines and halves the time of DreamGaussian [55] while retaining competitive performance.
Figure 4: Qualitative comparisons on mesh from 3DGS. MicroDreamer achieves an excellent efficiency-performance balance in terms of the visual quality of the mesh.
Quantitative comparisons.
We compare MicroDreamer with seven competitive baselines including Point-E [37], Shap-E [14], One-2-3-45 [23], TriplanceGaussian [67], Wonder3D [27], LGM [54] and DreamGaussian [55]. We record the optimization time  [^3] The detailed time recording method is provided in Appendix C. for each model on a single NVIDIA A100 (80GB) GPU, as shown in Fig. 5(a), and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works[23, 25, 55], as shown in Fig. 5(b). As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity. We provide detailed tables in Appendix E.
Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves. See detailed limitations and future work in Sec. 7.
(a) Optimization time for mesh on a single A100.
(b) CLIP similarity versus inference time.
Figure 5: Quantitative comparisons: efficiency versus quality. MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of the CLIP similarity [41]. The optimization-based methods are highlighted in red in (b).
5.3 Ablation Study
(a) Optimization space [52, 24].
(b) Optimization space [58].
(c) Iterative reconstruction.
(d) Number of camera views.
(e) Time schedule.
(f) Forward process.
Figure 6: Ablation study. (a-b) Optimization in pixel space accelerates the generation process. (c) Iterative reconstruction is necessary. (d) More camera views benefit. (e) A linear schedule is sufficient. (f) A hybrid forward process can be effective.
Ablation on optimization space.
From the results in Fig. 6(a) and Fig. 6(b), we can see that SIR-pixel achieves 2-3 times more acceleration than SIR-latent, verifying the benefits of SIR on enabling optimization in the pixel space in Sec. 3.3.
Ablation on iterative reconstruction.
We highlight the necessity of iterative reconstruction in Fig. 6(c). The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise. Although increasing $K$ improves the quality, we set $K\leq 30$ for efficiency.
Ablation on number of camera poses.
SIR necessitates more than one camera pose for reconstruction, as shown in Fig. 6(d). The performance increases with more camera poses. We choose $6$ on 3DGS with the Stable Zero123 [52] model to balance efficiency and quality. See Tab. 1 for the values in other settings.
Ablation on time schedule.
We compare our linear schedule, the random schedule employed in SDS [40], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$. As shown in Fig. 6(e), the schedules perform similarly. For simplicity and efficiency, we use the linear schedule by default.
Ablation on forward process.
Fig. 6(f) compares three forward processes on 3DGS with the Stable Zero123 [52] model. The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec. 3.2. See Tab. 1 for other settings.
6 Related work
Optimization-based 3D Generation.
Built upon text-to-image diffusion models, optimization-based approaches [40, 56, 19, 66, 5, 24, 46, 60, 55, 18, 53, 33, 62, 23, 22] generate 3D objects without additional 3D data. Among them, the most relevant line of work [40, 60, 18, 66, 56] proposes various distillation algorithms. Besides SDS [40], ProlificDreamer [60] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference. LucidDreamer [18] incorporates DDIM inversion to strengthen the forward process. As mentioned in Sec. 1, though promising, such methods suffer from inefficiency problems due to the large number of function evaluations (NFE) and optimization in the latent space. Distinct from existing algorithms, SIR mimics the 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion and enables optimization in the pixel space. Our experiments demonstrate that SIR is 5-20 times faster than SDS [40] on NeRF and twice faster than the most competitive baseline DreamGaussian [55] on meshes.
Feed-forward 3D Generation.
Contrary to optimization-based methods, feed-forward approaches[3, 4, 6, 14, 37, 26, 13, 35, 59, 61, 65, 67] elect to use large-scale 3D datasets[8, 7] training to achieve a process that can directly generate 3D objects. These methods are characterized by their speed but they often produce lower-quality 3D structures with simple textures. Several recent methods, utilizing more data, have exhibited possibilities. For instance, the large reconstruction model (LRM)[10, 13] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image. Instant3D[17] further expands on the LRM enabling it to accept multi-view input and reconstruct it into a NeRF. The recent work large multi-view Gaussian model (LGM)[54] employs a similar concept, generating 3DGS from multi-view images. Remarkably, our MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [41, 55].
Multi-view Prediction based 3D Generation.
There is also a line of work[27, 25] dedicated to enhancing the output of multi-view diffusion models to reconstruct 3D objects. Such systems, like SyncDreamer[25] and Wonder3D[27], train or fine-tune multi-view diffusion models on 3D datasets and use a single reconstruction process to obtain 3D objects from images without iteration. These methods typically rely on 3D prior knowledge and are limited to specific 3D representations [25, 27, 57] with a long reconstruction time. In contrast, thanks to the carefully designed iterative process in SIR, MicroDreamer applies to various 3D representations and is much more efficient (see Fig. 5(a) for a direct comparison).
7 Conclusion
We introduce SIR, an efficient and general algorithm mimicking the 3D reconstruction process to reduce the total NFEs in zero-shot 3D generation. Based on SIR, our method MicroDreamer achieves a 5 to 20 times speed increase for NeRF generation compared to SDS on various base models. Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches, with a competitive generation quality.
Limitations.
There is potential for further improving the efficiency of MicroDreamer via employing consistency models [50, 32] or alternative sampling models that require fewer steps [20, 63, 45]. Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion. Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
Potential Negative Impact.
As a generative approach, the paper can potentially be used for fabricating data and news. Moreover, the 3D content generated from input images may infringe on the privacy or copyright of others. While automatic detection might mitigate these issues, they warrant careful consideration.
8 Acknowledgement
Thanks Cheng Lu for the discussion when he was a PhD student at Tsinghua University.
Appendix 0.A Results of Applying SDS in Pixel Space.
We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq. 6. Fig. 7 shows the poor quality for single-step prediction from the diffusion model using Eq. 2. Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e. optimize 3D objects with 2D images like those generated in Fig. 7, the resulting 3D content would be of subpar quality, as the results presented in Fig. 8.
Figure 7: Results of single-step prediction. For each pair of images, the left is the input image and the right is the output of single-step prediction. The poor quality of the output makes it unsuitable for 3D optimization.
Figure 8: Results of applying SDS in pixel space. SDS optimizing in pixel space can not be effective.
Appendix 0.B Time Schedule of $t_{1}$
The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models. In our earlier exploration, we adopted $t_{1}=t^{2}_{2}/T$ for SIR based on MVDream [47] and ImageDream [58] on NeRF, and $t_{1}=0.6t_{2}$ for SIR based on Stable Zero123 [52, 24] on both NeRF and 3DGS. Notably, SIR based on ImageDream on 3DGS opted for $t_{1}=t_{2}$, i.e., adding noise for the forward process. As an advanced multi-view diffusion model, ImageDream can generate images with enhanced 3D consistency, obviating the need for inversion to maintain consistency. Fig. 9 shows some results of SIR for ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
Figure 9: Results of SIR based on ImageDream on 3DGS. These results images are rendered from 3DGS with no refinement stage.
Appendix 0.C More Implementation Details
We present the codebase, some details in metrics, and the data used in ablation sequentially.
0.C.1 Codebase
For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [9]. For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [55] [^4] https://github.com/dreamgaussian/dreamgaussian. The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^5] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k.
In Sec. 5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub, including Point-E [37] [^6] https://github.com/openai/point-e/, Shap-E [14] [^7] https://github.com/openai/shap-e, One-2-3-45 [23] [^8] https://github.com/One-2-3-45/One-2-3-45, TriplanceGaussian [67] [^9] https://github.com/VAST-AI-Research/TriplaneGaussian, Wonder3D [27] [^10] https://github.com/xxlong0/Wonder3D, LGM [54] [^11] https://github.com/3DTopia/LGM and DreamGaussian [55] [^12] https://github.com/dreamgaussian/dreamgaussian.
0.C.2 Details in Metric
All the results presented in Sec. 5.2 were on the mesh representation. As TriplaneGaussian [67] has no official mesh export code, we apply the mesh exported code from LGM [54] for it.
For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [55], the time of generating 3D from multi-view images for Wonder3D[27], and the time for meshing through a NeRF for LGM[54], respectively.
0.C.3 Data Used in Ablation Study
We chose seven images collected from the test dataset to perform our ablation study. We will include these images in the submitted supplementary material.
Appendix 0.D More Results on NeRF
In Fig. 10, we present a comparison of the generated quality between SIR and SDS on NeRF in the same short time under different base diffusion models [47, 52, 24, 58]. SIR can generate NeRF of high visual quality in a shorter time than SDS, as the results represented. Notably, SIR accelerates the generation process in ImageDream more, as it’s an advanced multi-view diffusion model that can produce higher-quality multi-view images and support fewer discretization steps for the diffusion process, lowering NFEs.
Figure 10: Comparison of the generated quality between SIR and SDS in the same short time on NeRF. SIR can generate NeRF of high visual quality in less time than SDS.
Appendix 0.E Quantative Comparison on 3DGS
We provide the detailed quantitative comparison table in Tab. 2. For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iter. As shown in Tab. 2, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
Table 2: Quantitative Comparisons. In terms of CLIP similarity, MicroDreamer achieves an excellent balance between efficiency and effectiveness.
Method CLIP-Similarity $\uparrow$ Generation time $\downarrow$ Extra refine $\downarrow$
Point-E 0.510 $\sim 24$s -
Shap-E 0.616 $\sim 5$s -
One-2-3-45 0.617 $\sim 42$s -
Wonder3D 0.699 $\sim 30$s $\sim 140$s
TriplaneGaussian 0.647 $\sim 7$s -
LGM 0.716 $\sim 3$s $\sim 50$s
DreamGaussian 0.698 $\sim 30$s $\sim 10$s
DreamGaussian-300 iter 0.656 $\sim 18$s $\sim 10$s
MicroDreamer-20 iter (Ours) 0.661 $\sim 18$s $\sim 4$s
MicroDreamer-30 iter (Ours) 0.687 $\sim 26$s $\sim 4$s

### v2

@@
 Abstract
-*Equal Contribution.
-Corresponding author.
 Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample.
-In this paper, we introduce score-based iterative reconstruction (SIR), an efficient and general algorithm for 3D generation with a multi-view score-based diffusion model.
-Given the images produced by the diffusion model, SIR reduces NFEs by repeatedly optimizing 3D parameters, unlike the single optimization in SDS, mimicking the 3D reconstruction process.
-With other improvements including optimization in the pixel space, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks.
-In particular, retaining a comparable performance, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field and takes about 20 seconds to generate meshes from 3D Gaussian splitting on a single A100 GPU, halving the time of the fastest zero-shot baseline, DreamGaussian.
+In this paper, we introduce score-based iterative reconstruction (SIR), an efficient and general algorithm mimicking a differentiable 3D reconstruction process to reduce the NFEs.
+Given a single set of images sampled from a multi-view score-based diffusion model, SIR repeatedly optimizes 3D parameters, unlike the single-step optimization in SDS.
+With other improvements in training, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks.
+In particular, retaining a comparable performance, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field and takes about 20 seconds to generate meshes from 3D Gaussian splatting on a single A100 GPU, halving the time of the fastest zero-shot baseline, DreamGaussian.
 Our code is available at https://github.com/ML-GSAI/MicroDreamer.
-Keywords:
-3D Generation Diffusion Model Multi-view Diffusion Score Distillation Sampling Zero-shot
 Figure 1: MicroDreamer can generate high-quality meshes in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data.
@@
 This advantage is particularly significant given that 3D data are costly and scarce.
-Despite their promising capabilities, optimization-based approaches suffer from low efficiency.
-Intrinsically, the inefficiency lies in the extensive number of function evaluations (NFEs), i.e.
+Despite their promising capabilities, optimization-based approaches suffer from low efficiency due to the extensive number of function evaluations (NFEs), i.e.
 forward passes of the diffusion model, required for each 3D object generation.
-Besides, within the commonly employed latent diffusion model (LDM) [43], the gradients of the 3D parameters need to be backpropagated through an encoder that maps rendered images to the latent space, further exacerbating the inefficiency of these approaches.
-To address the computational issues, this paper introduces an efficient and general algorithm termed score-based iterative reconstruction (SIR) for zero-shot 3D generation in Sec.
+In comparison, the multi-step reconstruction process of 3D representations that enable differentiable rendering, such as neural radiance field (NeRF) [35, 36] and 3D Gaussian splatting (3DGS) [37], produce 3D contents extremely fast because they do not involve large generative neural networks.
+However, such approaches rely on true 3D data (i.e.
+abundant real multi-view images like 100 in NeRF [35]), making it unfeasible to text-to-3D and image-to-3D generation tasks.
+There exist works [38, 39, 40, 41] for 3D generation attempt to generate multi-views first to reconstruct 3D object directly, while such methods may suffer from texture ambiguity [39] in generated views and are limited to specific 3D representations with a long reconstruction time (see Sec.
+5 for comparison).
+This paper introduces an efficient and general algorithm termed score-based iterative reconstruction (SIR) for zero-shot 3D generation inspiring from reconstruction and iterative optimization to address the computational issues in Sec.
 3.
 Similar to SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data.
-However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a single set of images produced by the diffusion model, mimicking the 3D reconstruction process to reduce the total NFEs (see Fig.
-3(a)).
-To obtain 3D-consistent and high-quality images as the ground truth for 3D reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration.
-Besides, even mapped back to the pixel space through the decoder in LDM, the refined images are still of high quality for reconstruction, enabling optimization in the pixel space to further speed up (see Fig.
-6(a)-6(b)).
+However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a set of images produced by the diffusion model, mimicking the efficient 3D reconstruction process to reduce the total NFEs (see Fig.
+4).
+To obtain 3D-consistent and high-quality images as the ground truth for better reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration.
 We delve into critical components beyond the algorithm to ensure compatibility with SIR.
@@
 As detailed in Sec.
-4, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3D Gaussian splatting (3DGS)[15].
-SIR and MicroDreamer are broadly applicable to both neural radiance fields (NeRF)[34] and 3DGS, and to both text-to-3D and image-to-3D tasks, as detailed in Sec.
+4, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3DGS.
+SIR and MicroDreamer are broadly applicable to both NeRF and 3DGS and both text-to-3D and image-to-3D tasks, as detailed in Sec.
 5.
@@
 Remarkably, MicroDreamer is on par with speed compared to feed-forward methods trained on extensive 3D data, with a very competitive CLIP similarity [45, 9].
-Figure 2: Overview of SIR and SDS.
-$\mathcal{E}$, $\mathcal{D}$ and $\epsilon_{\phi}$ are the encoder, decoder, and diffusion model in LDM [43].
-The solid and dashed arrows represent forward and backward passes respectively.
-The computation bottlenecks are highlighted in red.
-SIR is efficient because it lowers the NFEs in total and enables optimization in the pixel space.
+Figure 2: Overview of SIR.
+SIR is a zero-shot 3D generation method that marries the strengths of reconstruction and iterative optimization.
+SIR reutilizes the samples from diffusion multiple times through reconstruction, reducing the total NFEs and improving efficiency.
 2 Background
@@
 We denote the corresponding tunable parameters in both representations as $\theta$.
-Given camera poses $c$, both approaches define a differential rendering process, denoted by $g(\theta,c)$.
+Given camera poses $c$, both approaches define a differentiable rendering process, denoted by $g(\theta,c)$.
 They are proven efficient and effective in 3D reconstruction [35, 37] and generation [1, 9].
@@
 (1) but it can maintain the unique feature of an input image and reconstruct it accurately through the corresponding DDIM sampler in Eq.
-(3).
+(3) (see Fig.
+3a).
 The (one-step) DDIM inversion [52] is formulated as follows:
@@
 Among them, MVDream [42] takes text inputs and outputs multi-view images consistent in 3D.
-In contrast, Zero-1-to-3 [24] and ImageDream [58] focus on the Image-to-3D task, taking an additional reference image and its viewpoint as input.
+In contrast, Zero-1-to-3 [6] and ImageDream [44] focus on the Image-to-3D task, taking an additional reference image as input.
 These models output new viewpoint images consistent with the reference image.
@@
 For more details, please refer to the original papers and the source code in our supplementary materials.
-Notably, such multi-view diffusion models cannot provide a sufficient number of consistent multi-view images for classical 3D reconstruction, as shown in Fig.
-6(c).
+Notably, such multi-view diffusion models cannot provide a sufficient number of consistent multi-view images for 3D reconstruction directly, as shown in Fig.
+7(a).
+Given that the multi-view diffusion in [38, 39, 40, 41] often incorporate with additional 3D priors and do not align with our setting in Sec.
+4, we choose not to employ them in this paper.
 2.4 Optimization-based algorithms for 3D generation
@@
 3 Score-based iterative reconstruction
-We present the general framework of score-based iterative reconstruction (SIR), an efficient and versatile algorithm inspired by classical 3D reconstruction in Sec.
+We present the general framework of score-based iterative reconstruction (SIR), an efficient and versatile algorithm combining both the advantages of differentiable 3D reconstruction and optimization methods in Sec.
 3.1.
-SIR is designed to minimize the number of function evaluations (NFEs) typically required by existing methods.
-In Sec.
-3.2, we introduce a diffusion-based process to produce multi-view images as ground truth for reconstruction.
-In Sec.
-3.3, we demonstrate how SIR facilitates the optimization of 3D content directly within pixel space, thereby further boosting efficiency.
+SIR capitalizes on the efficiency of 3D reconstruction and aims to minimize the number of function evaluations (NFEs) typically required by existing methods while ensuring comparable quality through iterations.
+For better reconstruction, we introduce a carefully designed diffusion-based process to produce refined multi-view images as ground truth in Sec.
+3.2.
 3.1 Mimicking 3D reconstruction to reduce the NFEs
@@
 (5) for SDS as an example).
-Consequently, these methods necessitate a high number of function evaluations (NFEs), presenting a critical challenge to computational efficiency.
-Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of traditional 3D reconstruction—could substantially decrease the overall NFEs required.
-However, traditional 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [24, 47, 58, 52] (see Sec.
+Consequently, these methods necessitate high NFEs, presenting a critical challenge to computational efficiency.
+Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of differentiable 3D reconstruction could substantially decrease the overall NFEs required.
+However, 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [6, 42, 44, 43] (see Sec.
 6 for a review and discussion).
-Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple rounds of refinement, dubbed score-based iterative reconstruction (SIR), detailed as follows.
+Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple iterations of optimizing dubbed score-based iterative reconstruction (SIR), detailed as follows.
 Formally, SIR consists of $K$ reconstruction iterations.
-In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given an initial 3D parameter $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec.
-4) and employ a rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:
+In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given initial 3D parameters $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec.
+4) and employ rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:
 $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)}).$ (7)
@@
 3.2, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [42, 6, 44, 43].
-The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for the current iteration.
-In particular, starting from $\theta_{0}^{(k)}$, we optimize the 3D parameters w.r.t.
+The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for reconstruction in the current iteration.
+In particular, starting from $\theta_{0}^{(k)}$, keeping the camera poses $c^{(k)}$ unchanged, we optimize the 3D parameters w.r.t.
 the following reconstruction loss for $I^{(k)}$ steps:
@@
 Here we assume the rendered images and those sampled by the diffusion model are the same dimensional for simplicity.
-We will discuss how to deal with the latent diffusion model (LDM) [43] in Sec.
-3.3.
-Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ in Sec.
+In Appendix B.1 we will further discuss how to deal with the latent diffusion model (LDM) [17].
+Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ for reconstruction in Sec.
 3.2.
-3.2 Obtaining Multi-view Images for Reconstruction by Diffusion
-We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process.
+3.2 Refining multi-view images for reconstruction by diffusion
+We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process (see Fig. 3b).
+Figure 3: Overview of DDIM reconstruction and our refinement.
+(a) The inversion process can preserve the 3D consistency of input images, and the input images can be recovered through the corresponding DDIM sampler.
+(b) After the carefully designed refinement of $x$ by Eq.
+(9) and Eq.
+(10), the resulting $\hat{x}$ possesses enough high quality for 3D reconstruction.
 The diffusion model has two equivalent forward processes [16]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq.
@@
 In comparison, the noise-adding process is more efficient without function evaluation, and it aids in eliminating low-quality details through high-frequency noise.
-Conversely, the inversion process can preserve the 3D consistency and overall information of the current 3D object.
-Observations from existing experiments [38] indicate that diffusion models tend to generate details closer to time 00 and overall structure as they approach time $T$.
+Conversely, the inversion process can preserve the 3D consistency and overall information of the current 3D object (see Fig.
+3a).
+In addition, observations from existing experiments  [55] indicate that diffusion models tend to generate details closer to time 00 and overall structure as they approach time $T$.
 Inspired by this, we design a hybrid forward process that initially adds noise to time $t_{1}^{(k)}$ and then perform DDIM inversion to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise..
@@
 The optimal value of $\eta$ depends on the base model and we search for the best one in $\{0,0.5,1\}$ as detailed in Tab.
-1.
-The whole process is presented in Algorithm 1.
-Compared with existing methods such as SDS, SIR leverages several NFEs for multiple updates of 3D parameters and lowers the NFEs in total (see Fig. 3(a)).
+1 in Appendix A.
+The whole process of SIR is presented in Algorithm 1.
+Compared with existing generation methods such as SDS, SIR benefits from reconstruction, i.e.
+leverages several NFEs for multiple updates of 3D parameters, and lowers the NFEs in total (see Fig.
+4).
 1
@@
 Algorithm 1 Score-based iterative reconstruction
-3.3 Enabling Optimization in the Pixel Space
-Within the widely adopted framework of LDM [43], the rendered images are mapped through an encoder $\mathcal{E}$ [16] to a latent space, where the loss function is calculated.
-Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation.
-In such a case, the SIR loss in the latent space is given by:
-$\mathcal{J}_{\textrm{SIR-latent}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|\mathcal{E}(g(\theta,c^{(k)}))-\hat{x}^{(k)}\|.$ (11)
-An alternative approach maps the diffusion output back to the pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space.
-This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
-However, we argue that this is not feasible for existing optimization methods [40, 56, 60] such as SDS.
-Even reparameterized as the data predicting form in Eq.
-(6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix A).
-In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space.
-The SIR loss in the pixel space is formalized as:
-$\mathcal{J}_{\textrm{SIR-pixel}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|g(\theta,c^{(k)})-\mathcal{D}(\hat{x}^{(k)})\|,$ (12)
-which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as shown in Fig. 6(a)-6(b). It is set as the default loss throughout the paper.
 4 MicroDreamer
-In addition to our algorithm SIR, we carefully analyze and design other critical elements in the diffusion model and 3D training.
-Our comprehensive system is called MicroDreamer to highlight its efficiency.
+In addition to our algorithm SIR, we carefully analyze and design other critical elements in the 3D training and diffusion. Our comprehensive system is called MicroDreamer to highlight its efficiency.
 3D initialization and camera views.
-We utilize a direct reconstruction approach to initialize the 3D assets.
+We utilize a direct reconstruction approach to initialize the 3D content.
 Specifically, we optimize the loss function in Eq.
 (8) by several steps (see Tab.
-1 for detailed values) to update the 3D content, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion model.
+1 in Appendix A for detailed values) to update the 3D parameter, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion model.
 The camera views are uniformly sampled following the corresponding baselines [9, 56] except that the azimuth angles of different views in the same iteration are evenly distributed.
+In this setting, we are unable to directly utilize the diffusion model in [38, 39, 40] due to their fixed camera views conditioned generation.
 Annealed time schedule.
@@
 Our preliminary experiments in Fig.
-6(e) suggest that a linearly annealed schedule is sufficient.
+7(c) suggest that a linearly annealed schedule is sufficient.
 The endpoints of the schedule depend on the diffusion model, detailed in Tab.
-1.
-See Appendix B for the schedule of $t_{1}$.
+1 in Appendix A.
+See Appendix C for the schedule of $t_{1}$.
 Reference image loss for image-to-3D.
@@
 3DGS settings and mesh refinement.
-For simplicity, our approach to 3DGS largely adheres to the settings in DreamGaussian[55] unless specified.
-We incorporate a densify and prune procedure at every 100 updates during the initial 300 updates.
+For simplicity, our approach on 3DGS largely adheres to the settings in DreamGaussian [9] unless specified.
+We incorporate a densify and prune procedure at every 50 updates during the initial 200 updates.
 We follow the mesh extraction method in DreamGaussian  [9] and employ a threshold value of $0.2$ in the marching cube algorithm [57, 9].
-We utilize SIR to optimize the mesh texture in 3 iterations, with 15 steps per iteration.
-We use a noise-adding process and DDIM with $\eta=0$ for simplicity.
+We utilize SIR to optimize the exported mesh texture in 3 iterations, with 15 steps per iteration, and we use a noise-adding process and DDIM with $\eta=0$ for simplicity.
+5 Experiment
+We present the results on NeRF, 3DGS, and ablation sequentially.
+For a fair comparison, we utilize the widely used CLIP similarity [45] for quantitative comparison following DreamGaussian [9].
+We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0.
+For all methods, the corresponding images are rendered from NeRF in Sec.
+5.1 and mesh in Sec.
+5.2.
+We report the average CLIP similarities between these images and the reference image or text.
+See Appendix D for details of the checkpoint and codebase.
+5.1 Results on NeRF
+We apply SIR algorithm on NeRF [35, 36, 56] leveraging three multi-view diffusion model from MVDream [42], Stable Zero123 [43, 6], and ImageDream [44].
+For each model, we selected 6 input prompts from a widely used codebase [56] for testing, calculated NFEs, and recorded the change of the average CLIP similarity during the generation process.
+Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig.
+4.
+Besides, SIR can generate NeRF of high visual quality, as detailed shown in Appendix E, and the conclusion agrees with the results in Fig.
+4.
+(a) Comparison on MVDream
+(b) Comparison on Stable Zero123
+(c) Comparison on ImageDream
+Figure 4: Comparison of SIR and SDS on NeRF. The change of the CLIP similarity in the generation process and final NFEs on different models. SIR achieves a 5-20 times acceleration and lowers NFEs.
+5.2 Results on 3D Gaussian splatting
+Qualitative comparisons.
+We show the generated 3D meshes between the fastest optimization-based baseline DreamGaussian [9] and ours in Fig.
+5.
+MicroDreamer halves the generation time of DreamGaussian while retaining competitive performance.
+We provide more visualization results of our methods in Appendix E.
+Figure 5: Qualitative comparisons on mesh from 3DGS. MicroDreamer retains a comparable performance in terms of the visual quality of the mesh with higher efficiency.
+Quantitative comparisons.
+We compare MicroDreamer with seven competitive baselines including Point-E [24], Shap-E [23], One-2-3-45 [31], TriplanceGaussian [33], Wonder3D [39], LGM [30] and DreamGaussian [9].
+We record the optimization time  [^3] The detailed time recording method is provided in Appendix D.
+for each model on a single NVIDIA A100 (80GB) GPU, and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works [31, 38, 9], as shown in Fig.
+7.
+As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity.
+We provide the time record and detailed tables in Appendix F.
+Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves.
+See detailed limitations and future work in Sec.
+7 and Appendix E.
+Figure 6: Quantitative comparisons. Our MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of CLIP similarity. The zero-shot methods are highlighted in red.
+[…]
+(a) Iterative reconstruction.
+(b) Camera views.
+(c) Time schedule.
+(d) Forward process.
+Figure 7: Ablation study. (a) Iterative reconstruction is necessary. (b) More camera views benefit. (c) A linear schedule is sufficient. (d) A hybrid forward process can be effective.
+5.3 Ablation study
+Ablation on iterative reconstruction.
+Fig.
+7(a) shows the necessity of iterative reconstruction.
+The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise.
+Although increasing $K$ improves quality, we set $K\leq 30$ for efficiency.
+Ablation on number of camera poses.
+SIR necessitates more than 1 camera pose for reconstruction, as shown in Fig.
+7(b).
+The performance increases with more camera poses.
+We choose $6$ on 3DGS with Stable Zero123 [43] model to balance efficiency and quality.
+See Tab.
+1 for values in other settings.
+Ablation on time schedule.
+We compare our linear schedule, the random schedule employed in SDS [1], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$.
+As shown in Fig.
+7(c), the schedules perform similarly.
+For simplicity and efficiency, we use the linear schedule by default.
+Ablation on forward process.
+Fig.
+7(d) compares three forward processes on 3DGS with the Stable Zero123 [43] model.
+The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec.
+3.2.
+See Tab.
+1 in Appendix A for other settings.
+6 Related work
+Optimization-based 3D generation.
+Built upon text-to-image diffusion models, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 58, 59, 60] usually generate 3D objects without additional 3D data.
+Among them, the most relevant line of work [1, 8, 10, 4, 2] proposes various distillation algorithms.
+Besides SDS [1], ProlificDreamer [8] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference.
+LucidDreamer [10] incorporates DDIM inversion to strengthen the forward process.
+Though promising, such methods suffer from inefficiency problems due to large NFEs.
+Distinct from existing algorithms, SIR mimics 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion.
+Our experiments demonstrate that SIR is 5-20 times faster than SDS [1] on NeRF and twice as fast as the most competitive baseline DreamGaussian [9] on meshes.
+Feed-forward 3D generation.
+Contrary to optimization-based methods, feed-forward methods [20, 21, 22, 23, 24, 25, 26, 27, 28, 61, 29, 31, 32, 33, 34, 62, 63, 64, 65] elect to use large-scale 3D datasets [66, 67] training to achieve a process that can directly generate 3D objects.
+Early methods are characterized by their speed but often produce lower-quality 3D structures with simple textures.
+Several recent methods, utilizing more data, have exhibited possibilities.
+Large reconstruction model (LRM) [68, 26] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image.
+Instant3D [69] further expands on LRM enabling it to accept multi-view input and reconstruct it into a NeRF.
+Remarkably, MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [45, 9].
+Multi-view prediction based 3D generation.
+There is also a line of work [38, 39, 40, 41] dedicated to enhancing the output of multi-view diffusion models by training on 3D datasets to reconstruct 3D objects using a single reconstruction process without iteration.
+These methods like Wonder3D[39] typically rely on 3D prior knowledge and are limited to specific 3D representations with a long reconstruction time.
+In contrast, thanks to the iterative process in SIR, MicroDreamer applies to various 3D representations and is more efficient (see Fig.
+7 for a direct comparison).
+7 Conclusion
+We introduce SIR, an efficient and general algorithm combining both the strengths of 3D reconstruction and iterative optimization to reduce total NFEs in zero-shot 3D generation.
+SIR achieves a 5 to 20 times speed increase for NeRF generation compared to SDS.
+Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches with a competitive generation quality.
+Limitations and broader impact.
+There is potential for further improving MicroDreamer’s efficiency via employing consistency models [70, 71] or alternative sampling models that require fewer steps [72, 73].
+Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion.
+Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
+Also, MicroDreamer can provide convenience for artists to create 3D assets.
+However, as a generative approach, our method may also be used for fabricating data and news, which warrants careful consideration.
+8 Acknowledgement
+Thanks Cheng Lu for the discussion when he was a PhD student at Tsinghua University.
+Appendix A Key hyperparameters of MicroDreamer
 Table 1: Key hyperparameters of MicroDreamer on three base diffusion models.
@@
 Reference opacity loss - 0.001 (0.01) 0
-5 Experiment
-We present the results on NeRF, 3DGS, and ablation sequentially.
-For a fair comparison, we utilize the widely used CLIP similarity [41] for quantitative comparison following DreamGaussian [55].
-We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0.
-For all methods, the corresponding images are rendered from NeRF in Sec.
-5.1 and mesh in Sec.
-5.2.
-We report the average CLIP similarities between these images and the reference image or text.
-See Appendix C for details of the checkpoint and codebase.
-5.1 Results on NeRF
-We apply SIR algorithm on NeRF[34, 36, 9] leveraging three multi-view diffusion model from MVDream[47], Stable Zero123[52, 24], and ImageDream[58].
-For each model, we selected 6 input prompts from a widely used codebase [9] for testing, calculated NFEs, and recorded the change of the average CLIP similarity in the generation process.
-Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig.
-3.
-Besides, SIR can generate NeRF of high visual quality, detailed in Appendix D, and the conclusion agrees with the results in Fig.
-3.
-(a) Comparsion w.r.t. NFEs.
-(b) Comparison on MVDream[47]
-(c) Comparison on Stable Zero123[52, 24]
-(d) Comparison on ImageDream[58]
-Figure 3: Comparison of SIR and SDS on NeRF.
-(a) shows that SIR lowers the total NFEs.
-(b-d) show the change of the CLIP similarity in the generation process under different models and SIR achieves a 5-20 times acceleration approximately.
-5.2 Results on 3D Gaussian Splatting
-Qualitative comparisons.
-We show the generated 3D objects amongst various methods and ours in Fig.
-4.
-The baselines include representative feed-forward models One-2-3-45 [23] and TriplaneGaussian [67], and the fastest optimization-based model DreamGaussian [55].
-We present the exported mesh from these models.
-MicroDreamer generates higher quality mesh exported from 3DGS than the feed-forward baselines and halves the time of DreamGaussian [55] while retaining competitive performance.
-Figure 4: Qualitative comparisons on mesh from 3DGS. MicroDreamer achieves an excellent efficiency-performance balance in terms of the visual quality of the mesh.
-Quantitative comparisons.
-We compare MicroDreamer with seven competitive baselines including Point-E [37], Shap-E [14], One-2-3-45 [23], TriplanceGaussian [67], Wonder3D [27], LGM [54] and DreamGaussian [55].
-We record the optimization time  [^3] The detailed time recording method is provided in Appendix C.
-for each model on a single NVIDIA A100 (80GB) GPU, as shown in Fig.
-5(a), and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works[23, 25, 55], as shown in Fig.
-5(b).
-As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity.
-We provide detailed tables in Appendix E.
-Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves. See detailed limitations and future work in Sec. 7.
-(a) Optimization time for mesh on a single A100.
-(b) CLIP similarity versus inference time.
-Figure 5: Quantitative comparisons: efficiency versus quality.
-MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of the CLIP similarity [41].
-The optimization-based methods are highlighted in red in (b).
-5.3 Ablation Study
-(a) Optimization space [52, 24].
-(b) Optimization space [58].
-(c) Iterative reconstruction.
-(d) Number of camera views.
-(e) Time schedule.
-(f) Forward process.
-Figure 6: Ablation study.
-(a-b) Optimization in pixel space accelerates the generation process.
-(c) Iterative reconstruction is necessary.
-(d) More camera views benefit.
-(e) A linear schedule is sufficient.
-(f) A hybrid forward process can be effective.
-Ablation on optimization space.
-From the results in Fig.
-6(a) and Fig.
-6(b), we can see that SIR-pixel achieves 2-3 times more acceleration than SIR-latent, verifying the benefits of SIR on enabling optimization in the pixel space in Sec.
-3.3.
-Ablation on iterative reconstruction.
-We highlight the necessity of iterative reconstruction in Fig.
-6(c).
-The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise.
-Although increasing $K$ improves the quality, we set $K\leq 30$ for efficiency.
-Ablation on number of camera poses.
-SIR necessitates more than one camera pose for reconstruction, as shown in Fig.
-6(d).
-The performance increases with more camera poses.
-We choose $6$ on 3DGS with the Stable Zero123 [52] model to balance efficiency and quality.
-See Tab.
-1 for the values in other settings.
-Ablation on time schedule.
-We compare our linear schedule, the random schedule employed in SDS [40], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$.
-As shown in Fig.
-6(e), the schedules perform similarly.
-For simplicity and efficiency, we use the linear schedule by default.
-Ablation on forward process.
-Fig.
-6(f) compares three forward processes on 3DGS with the Stable Zero123 [52] model.
-The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec.
-3.2.
-See Tab.
-1 for other settings.
-6 Related work
-Optimization-based 3D Generation.
-Built upon text-to-image diffusion models, optimization-based approaches [40, 56, 19, 66, 5, 24, 46, 60, 55, 18, 53, 33, 62, 23, 22] generate 3D objects without additional 3D data.
-Among them, the most relevant line of work [40, 60, 18, 66, 56] proposes various distillation algorithms.
-Besides SDS [40], ProlificDreamer [60] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference.
-LucidDreamer [18] incorporates DDIM inversion to strengthen the forward process.
-As mentioned in Sec.
-1, though promising, such methods suffer from inefficiency problems due to the large number of function evaluations (NFE) and optimization in the latent space.
-Distinct from existing algorithms, SIR mimics the 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion and enables optimization in the pixel space.
-Our experiments demonstrate that SIR is 5-20 times faster than SDS [40] on NeRF and twice faster than the most competitive baseline DreamGaussian [55] on meshes.
-Feed-forward 3D Generation.
-Contrary to optimization-based methods, feed-forward approaches[3, 4, 6, 14, 37, 26, 13, 35, 59, 61, 65, 67] elect to use large-scale 3D datasets[8, 7] training to achieve a process that can directly generate 3D objects.
-These methods are characterized by their speed but they often produce lower-quality 3D structures with simple textures.
-Several recent methods, utilizing more data, have exhibited possibilities.
-For instance, the large reconstruction model (LRM)[10, 13] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image.
-Instant3D[17] further expands on the LRM enabling it to accept multi-view input and reconstruct it into a NeRF.
-The recent work large multi-view Gaussian model (LGM)[54] employs a similar concept, generating 3DGS from multi-view images.
-Remarkably, our MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [41, 55].
-Multi-view Prediction based 3D Generation.
-There is also a line of work[27, 25] dedicated to enhancing the output of multi-view diffusion models to reconstruct 3D objects.
-Such systems, like SyncDreamer[25] and Wonder3D[27], train or fine-tune multi-view diffusion models on 3D datasets and use a single reconstruction process to obtain 3D objects from images without iteration.
-These methods typically rely on 3D prior knowledge and are limited to specific 3D representations [25, 27, 57] with a long reconstruction time.
-In contrast, thanks to the carefully designed iterative process in SIR, MicroDreamer applies to various 3D representations and is much more efficient (see Fig.
-5(a) for a direct comparison).
-7 Conclusion
-We introduce SIR, an efficient and general algorithm mimicking the 3D reconstruction process to reduce the total NFEs in zero-shot 3D generation.
-Based on SIR, our method MicroDreamer achieves a 5 to 20 times speed increase for NeRF generation compared to SDS on various base models.
-Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches, with a competitive generation quality.
-Limitations.
-There is potential for further improving the efficiency of MicroDreamer via employing consistency models [50, 32] or alternative sampling models that require fewer steps [20, 63, 45].
-Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion.
-Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
-Potential Negative Impact.
-As a generative approach, the paper can potentially be used for fabricating data and news.
-Moreover, the 3D content generated from input images may infringe on the privacy or copyright of others.
-While automatic detection might mitigate these issues, they warrant careful consideration.
-8 Acknowledgement
-Thanks Cheng Lu for the discussion when he was a PhD student at Tsinghua University.
-Appendix 0.A Results of Applying SDS in Pixel Space.
+Appendix B The impact of latent diffusion model on 3D generation
+Instead of applying the diffusion model directly in the pixel space to generate images, existing research [17] tends to use a VAE [75, 76] encoder to compress high-dimensional image data into a low-dimensional latent space first.
+Then a diffusion process is employed within this latent space to generate new images, which are finally decoded back by a VAE decoder to the pixel space to achieve high-quality images.
+We refer to this type of diffusion model as the latent diffusion model (LDM).
+The framework of LDM is adopted widely [6, 42, 44, 43], which leads to a significant impact on 3D generation.
+B.1 SIR enabling optimization in the pixel space
+Within the framework of LDM [17], the rendered images are mapped through an encoder $\mathcal{E}$ [75, 76] to a latent space, where the loss function is calculated.
+Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation.
+In such a case, the SIR loss in the latent space is given by:
+$\mathcal{J}_{\textrm{SIR-latent}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|\mathcal{E}(g(\theta,c^{(k)}))-\hat{x}^{(k)}\|.$ (11)
+An alternative approach maps the diffusion output back to the pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space.
+This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
+However, we argue that this is not feasible for existing optimization methods [1, 2, 8] such as SDS.
+Even reparameterized as the data predicting form in Eq.
+(6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix B.2).
+In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space.
+The SIR loss in the pixel space is formalized as:
+$\mathcal{J}_{\textrm{SIR-pixel}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|g(\theta,c^{(k)})-\mathcal{D}(\hat{x}^{(k)})\|,$ (12)
+which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as the ablation results shown in Fig. 8. It is set as the default loss throughout the paper.
+Figure 8: Optimization in pixel space accelerates generation.
+B.2 Faliure of applying SDS in pixel space.
 We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq.
-6.
+(6).
 Fig.
 9 shows the poor quality for single-step prediction from the diffusion model using Eq.
-2.
+(2).
 Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e.
-optimize 3D objects with 2D images like those generated in Fig.
-7, the resulting 3D content would be of subpar quality, as the results presented in Fig.
-8.
+optimizing 3D objects with 2D images like those generated in Fig.
+9, the resulting 3D content would be of subpar quality, as the results presented in Fig.
+10.
 Figure 9: Results of single-step prediction.
@@
 Figure 10: Results of applying SDS in pixel space. SDS optimizing in pixel space can not be effective.
-Appendix 0.B Time Schedule of $t_{1}$
+Appendix C Time schedule of $t_{1}$
 The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models.
@@
 Fig.
-9 shows some results of SIR for ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
+11 shows some results of SIR on ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
 Figure 11: Results of SIR based on ImageDream on 3DGS. These results images are rendered from 3DGS with no refinement stage.
-Appendix 0.C More Implementation Details
-We present the codebase, some details in metrics, and the data used in ablation sequentially.
-0.C.1 Codebase
+Appendix D More implementation details
+Codebase.
 For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [56].
-For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [55] [^4] https://github.com/dreamgaussian/dreamgaussian.
-The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [^5] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k.
+For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [9].
+The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [45, 77, 78].
 In Sec.
-5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub, including Point-E [37] [^6] https://github.com/openai/point-e/, Shap-E [14] [^7] https://github.com/openai/shap-e, One-2-3-45 [23] [^8] https://github.com/One-2-3-45/One-2-3-45, TriplanceGaussian [67] [^9] https://github.com/VAST-AI-Research/TriplaneGaussian, Wonder3D [27] [^10] https://github.com/xxlong0/Wonder3D, LGM [54] [^11] https://github.com/3DTopia/LGM and DreamGaussian [55] [^12] https://github.com/dreamgaussian/dreamgaussian.
-0.C.2 Details in Metric
-All the results presented in Sec. 5.2 were on the mesh representation. As TriplaneGaussian [67] has no official mesh export code, we apply the mesh exported code from LGM [54] for it.
+5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub.
+See Tab.
+2 for details about the URL and licenses of the assets we used in this paper.
+Table 2: Codebase. URL and licenses are provided for the open-source assets we used in this paper.
+Details in metric.
+All the results presented in Sec.
+5.2 were on the mesh representation.
+As TriplaneGaussian [33] has no official mesh export code, we apply the mesh exported code from LGM [30] for it.
 For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [9], the time of generating 3D from multi-view images for Wonder3D [39], and the time for meshing through a NeRF for LGM [30], respectively.
-0.C.3 Data Used in Ablation Study
-We chose seven images collected from the test dataset to perform our ablation study. We will include these images in the submitted supplementary material.
-Appendix 0.D More Results on NeRF
+Data used in ablation study.
+We chose seven images collected from the test dataset to perform our ablation study in Fig. 7 and Fig. 8. We will include these images in the submitted supplementary material.
+Safeguards.
+Our method, which requires no additional models and doesn’t release any new pre-trained model, can ensure security due to the inherent safe configuration of the employed 2D multi-view diffusion models [42, 43, 44].
+Appendix E More results
+Results on 3DGS.
+We provide more results on mesh exported from 3DGS for MicroDreamer-20 iters in Fig. 12.
+Figure 12: More results for MicroDreamer-20 iters.
+Results on NeRF.
 In Fig.
-10, we present a comparison of the generated quality between SIR and SDS on NeRF in the same short time under different base diffusion models [47, 52, 24, 58].
+13, we present a comparison of the generated quality between SIR and SDS on NeRF in the same short time under different base diffusion models [42, 43, 6, 44].
 SIR can generate NeRF of high visual quality in a shorter time than SDS, as the results represented.
@@
 Figure 13: Comparison of the generated quality between SIR and SDS in the same short time on NeRF. SIR can generate NeRF of high visual quality in less time than SDS.
-Appendix 0.E Quantative Comparison on 3DGS
+Limitations.
+In Fig.
+14, we show some of the less successful cases generated by our method.
+Due to the limited ability of the diffusion model, inputs with complex geometric structures such as central hollowing may not result in satisfactory mesh outcomes.
+Furthermore, existing diffusion models struggle to generate the rear view of an object, which often leads to poorly textured back surfaces of the generated meshes.
+These problems may be solved as the quality of the images the multi-view diffusion model generated improves.
+Figure 14: Limitations. Visualization of some less satisfactory cases generated by our methods.
+Appendix F Quantative comparison on 3DGS
 We provide the detailed quantitative comparison table in Tab.
-2.
-For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iter.
+3.
+For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iters.
 As shown in Tab.
-2, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
+3, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
 Table 3: Quantitative comparisons. In terms of CLIP similarity, MicroDreamer achieves an excellent balance between efficiency and effectiveness.
@@
 DreamGaussian 0.698 $\sim 30$s $\sim 10$s
-DreamGaussian-300 iter 0.656 $\sim 18$s $\sim 10$s
-MicroDreamer-20 iter (Ours) 0.661 $\sim 18$s $\sim 4$s
-MicroDreamer-30 iter (Ours) 0.687 $\sim 26$s $\sim 4$s
+DreamGaussian-300 iters 0.656 $\sim 18$s $\sim 10$s
+MicroDreamer-20 iters (Ours) 0.678 $\sim 18$s $\sim 4$s
+MicroDreamer-30 iters (Ours) 0.687 $\sim 26$s $\sim 4$s

### v3

@@
-MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction
+MicroDreamer: Efficient 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction
 Abstract
-Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample.
-In this paper, we introduce score-based iterative reconstruction (SIR), an efficient and general algorithm mimicking a differentiable 3D reconstruction process to reduce the NFEs.
+Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample and the limitation of optimization confined to latent space.
+This paper introduces score-based iterative reconstruction (SIR), an efficient and general algorithm mimicking a differentiable 3D reconstruction process to reduce the NFEs and enable optimization in pixel space.
 Given a single set of images sampled from a multi-view score-based diffusion model, SIR repeatedly optimizes 3D parameters, unlike the single-step optimization in SDS.
 With other improvements in training, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks.
-In particular, retaining a comparable performance, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field and takes about 20 seconds to generate meshes from 3D Gaussian splatting on a single A100 GPU, halving the time of the fastest zero-shot baseline, DreamGaussian.
+In particular, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field while retaining a comparable performance and takes about 20 seconds to create meshes from 3D Gaussian splatting on a single A100 GPU, halving the time of the fastest optimization-based baseline DreamGaussian with significantly superior performance compared to the measurement standard deviation.
 Our code is available at https://github.com/ML-GSAI/MicroDreamer.
-Figure 1: MicroDreamer can generate high-quality meshes in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data.
-1 Introduction
-Recently, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] particularly score distillation sampling (SDS) [1, 2] have emerged as promising avenues for 3D generation based on text-to-image diffusion models [14, 15, 16, 17, 18, 19].
-These approaches are appealing due to their minimal, or even zero reliance on 3D data, in contrast to the data-intensive requirements of feed-forward approaches [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
+Index Terms:
+3D Generation, Diffusion Model, Multi-view Diffusion, Score Distillation Sampling
+I Introduction
+Recently, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] particularly score distillation sampling (SDS) [1, 2] have emerged as promising avenues for 3D generation based on text-to-image diffusion models [20, 21, 22, 23, 24, 25].
+These approaches are appealing due to their minimal, or even zero reliance on 3D data, in contrast to the data-intensive requirements of feed-forward approaches [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53].
 This advantage is particularly significant given that 3D data are costly and scarce.
@@
 forward passes of the diffusion model, required for each 3D object generation.
-In comparison, the multi-step reconstruction process of 3D representations that enable differentiable rendering, such as neural radiance field (NeRF) [35, 36] and 3D Gaussian splatting (3DGS) [37], produce 3D contents extremely fast because they do not involve large generative neural networks.
-However, such approaches rely on true 3D data (i.e.
-abundant real multi-view images like 100 in NeRF [35]), making it unfeasible to text-to-3D and image-to-3D generation tasks.
-There exist works [38, 39, 40, 41] for 3D generation attempt to generate multi-views first to reconstruct 3D object directly, while such methods may suffer from texture ambiguity [39] in generated views and are limited to specific 3D representations with a long reconstruction time (see Sec.
-5 for comparison).
-This paper introduces an efficient and general algorithm termed score-based iterative reconstruction (SIR) for zero-shot 3D generation inspiring from reconstruction and iterative optimization to address the computational issues in Sec.
-3.
-Similar to SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data.
+Moreover, when adopting the latent diffusion model (LDM) [23] framework, most of these approaches can only compute loss in latent space rather than pixel space, which requires backpropagation through an encoder [54, 55] and further lowers the efficiency.
+Even if the loss function can be written in a data-predicting form [4], this type of loss is challenging to optimize effectively (see evidence in Fig.
+11).
+The fastest approach DreamGaussian [8] still requires about 40 seconds to generate a 3D object even employing 3D Gaussian splatting [56].
+Figure 1: MicroDreamer surpasses the fastest optimization-based baseline DreamGaussian [8] in terms of both efficiency and sample quality.
+The optimization-based methods are highlighted in red.
+See Tab.
+III for a comprehensive comparison with more baselines.
+Figure 2: MicroDreamer can generate a high-quality mesh (as illustrated above) in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data.
+See supplementary materials for 3D visualization.
+In comparison, the multi-step reconstruction process of 3D representations that enable differentiable rendering, such as neural radiance field (NeRF) [57, 58] and 3D Gaussian splatting (3DGS) [56], produces 3D contents extremely fast because they do not involve large generative neural networks.
+However, such approaches rely on true 3D data, i.e.
+abundant real multi-view images, making them unfeasible for text-to-3D and image-to-3D generation tasks.
+There exist works [59, 60, 61, 62, 63, 64, 65, 66] for 3D generation attempt to generate multi-views first to reconstruct 3D object directly.
+Still, such methods may require the diffusion model to simultaneously generate multi-view images with their corresponding 3D priors [60, 63], and they require longer than one minute to reconstruct a 3D object.
+(see Sec.
+II for a review and Sec.
+VI for comparison).
+To speed up the generation process, this paper presents an efficient and general 3D generation algorithm termed score-based iterative reconstruction (SIR), leveraging reconstruction to reduce NFEs and enable optimization in pixel space.
+Like SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data or 3D prior.
 However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a set of images produced by the diffusion model, mimicking the efficient 3D reconstruction process to reduce the total NFEs (see Fig.
-4).
+7).
 To obtain 3D-consistent and high-quality images as the ground truth for better reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration.
-We delve into critical components beyond the algorithm to ensure compatibility with SIR.
-The comprehensive system is named by MicroDreamer, highlighting its efficiency for 3D generation.
+Besides, even mapped back to pixel space through the decoder in LDM, the refined images are still of high quality for reconstruction, enabling optimization in pixel space to speed up further (see Fig.
+10).
+We provide a general and compatible configuration for SIR, and the comprehensive system is named MicroDreamer, highlighting its efficiency for 3D generation.
 As detailed in Sec.
-4, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3DGS.
-SIR and MicroDreamer are broadly applicable to both NeRF and 3DGS and both text-to-3D and image-to-3D tasks, as detailed in Sec.
-5.
+V, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3DGS.
+Comprehensive studies demonstrate the generality and efficiency of our proposed method.
+In particular, SIR and MicroDreamer broadly apply to NeRF and 3DGS and both text-to-3D and image-to-3D tasks, as detailed in Sec.
+VI.
 Employing three widely adopted multi-view diffusion models [67, 68, 69], we systematically compare SIR and SDS for NeRF generation.
 Retaining a comparable performance, SIR can accelerate the generation process by 5 to 20 times.
-Besides, MicroDreamer can efficiently produce 3DGS and further refine it into high-quality meshes, delivering consistent 3D meshes in about 20 seconds on a single A100 GPU – about twice as fast as the most competitive optimization-based alternatives, DreamGaussian [9].
-Remarkably, MicroDreamer is on par with speed compared to feed-forward methods trained on extensive 3D data, with a very competitive CLIP similarity [45, 9].
+Besides, MicroDreamer can efficiently produce 3DGS and further refine it into high-quality meshes, delivering consistent 3D meshes in about 20 seconds on a single A100 GPU – about twice as fast as the most competitive optimization-based alternatives, DreamGaussian [8], with significantly (compared to the measurement standard deviation) superior performance.
+Remarkably, MicroDreamer is on par with speed compared to representative feed-forward methods [70] trained on extensive 3D data, with a very competitive CLIP similarity [71].
 Figure 3: Overview of SIR.
-SIR is a zero-shot 3D generation method that marries the strengths of reconstruction and iterative optimization.
-SIR reutilizes the samples from diffusion multiple times through reconstruction, reducing the total NFEs and improving efficiency.
-2 Background
-We present background on 3D representations, diffusion models, multi-view diffusion models, and current optimization-based 3D generation methods.
-2.1 3D representation
+SIR is an optimization-based 3D generation method that marries the strengths of reconstruction and iterative optimization.
+SIR reutilizes the samples from diffusion multiple times through reconstruction, reducing the total NFEs, enabling optimization in pixel space, and improving efficiency.
+II Related work
+Optimization-based 3D generation.
+Built upon text-to-image diffusion models, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] usually generate 3D objects without additional 3D data.
+Among them, the most relevant line of work [1, 7, 9, 4, 2, 17, 18, 19] proposes various distillation algorithms.
+Besides SDS [1], ProlificDreamer [7] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference.
+LucidDreamer [9] incorporates DDIM inversion to strengthen the forward process.
+ReconFusion [15] utilizes diffusion priors to generate novel views for single-step optimization in sparse view reconstruction tasks, aiming to enhance the quality of scene reconstruction.
+Though promising, many of these methods suffer from inefficiency problems due to large NFEs and optimization in latent space.
+Distinct from existing algorithms, SIR mimics 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion.
+Our experiments demonstrate that SIR is 5-20 times faster than SDS [1] on NeRF and twice as fast as the most competitive baseline DreamGaussian [8] with better generation quality on meshes.
+Feed-forward 3D generation.
+Contrary to optimization-based methods, feed-forward methods [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] use large-scale 3D datasets [72, 73] training to achieve a process that can directly generate 3D objects.
+Early methods are characterized by their speed but often produce lower-quality 3D structures with simple textures.
+Several recent methods [70, 32, 36, 37, 48, 53] have exhibited the feasibility of training a Transformer utilizing more data to achieve reliable 3D content from a single image.
+Generating 3D content from a single view can be quite challenging.
+Therefore, several follow-up works [38, 39, 40, 49, 51, 42, 43, 45] have leveraged multi-view diffusion models to generate multiple views as input and then trained a feed-forward model to produce 3D content from these views.
+Remarkably, MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [71].
+Multi-view prediction based 3D generation.
+There is also a line of work [59, 60, 61, 62, 63, 64, 65, 66] dedicated to enhancing the output of multi-view diffusion models by training on 3D datasets to reconstruct 3D objects using a single reconstruction process with no or few iterations.
+Some approaches like Wonder3D [60] typically rely on 3D prior knowledge and are limited to specific 3D representations with a long reconstruction time.
+Another aspect of works like IM-3D [64] and V3D [65] involves fine-tuning the video diffusion model using 3D data and employing the generated 3D-aware multi-view images as ground truth for reconstruction.
+The efficiency of these methods is limited by the sampling efficiency of the video diffusion model and the long reconstruction time.
+Consequently, the total time required to generate a 3D object using these methods typically exceeds one minute.
+In contrast, thanks to the carefully designed iterative process in SIR, MicroDreamer is more efficient and applies to various 3D representations (see Tab.
+III for comparison).
+III Background
+We present background on 3D representations, diffusion models, multi-view diffusion models, and current optimization-based 3D generation methods sequentially.
+III-A 3D representation
 Neural radiance fields [57, 58] (NeRF) and 3D Gaussian splatting [56] (3DGS) have emerged as popular 3D representations.
 NeRF employs an MLP to predict the colors and density of the input space coordinates.
-3DGS consists of multiple 3D Gaussians parameterized by the colors, center, scale, and rotation quaternion.
+3DGS consists of multiple 3D Gaussians parameterized by the colors, centers, scales, and rotation quaternions.
 We denote the corresponding tunable parameters in both representations as $\theta$.
@@
 They are proven efficient and effective in 3D reconstruction [57, 56] and generation [1, 8].
-2.2 Diffusion models
+III-B Diffusion models
 A diffusion model [20, 21, 22] consists of a forward process and a sampling process.
@@
 For any $t\in(0,T)$, the noise-adding process can be written as follows:
-$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon:=\textrm{Noise-adding}(x_{0},0\rightarrow t),\quad\epsilon\in\mathcal{N}(0,I),$ (1)
+$x_{t}$ $=\alpha_{t}x_{0}+\sigma_{t}\epsilon$
+[…]
 where the coefficients $\alpha_{t}$ and $\sigma_{t}$ form the noise schedule of the diffusion model.
@@
 Among existing samplers [74, 75, 76, 77, 78, 79], the denoising diffusion implicit models (DDIM) [80] facilitate a sequence of samplers with random noise control $\eta$.
-When $\eta=0$, it solves the equivalent probability flow ordinary differential equation (ODE) [16] of the diffusion model and enjoys a fast sampling process with fewer number of function evaluations (NFEs) [^1] Throughout the paper, we refer to the number of forward passes through $\epsilon_{\phi}$ as NFEs..
+When $\eta=0$, it solves the equivalent probability flow ordinary differential equation (ODE) [22] of the diffusion model and enjoys a fast sampling process with a small number of function evaluations (NFEs) [^1] Throughout the paper, we refer to the number of forward passes through $\epsilon_{\phi}$ as NFEs..
 In this setting, the one-step sampling of DDIM is given by:
-$x_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}+\left(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Sampler}}(x_{t},t\rightarrow t-1).$ (3)
+$x_{t-1}$ $=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}+\left(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t)$
+$:={\textrm{Sampler}}(x_{t},t\rightarrow t-1).$ (3)
 We refer the readers to the original paper [80] for the formula of $\eta>0$.
@@
 (3) (see Fig.
-3a).
+5a).
 The (one-step) DDIM inversion [80] is formulated as follows:
-$x_{t+1}=\frac{\alpha_{t+1}}{\alpha_{t}}x_{t}+\left(\sigma_{t+1}-\frac{\alpha_{t+1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Inversion}}(x_{t},t\rightarrow t+1).$ (4)
-2.3 Multi-view diffusion models
+$x_{t+1}$ $=\frac{\alpha_{t+1}}{\alpha_{t}}x_{t}+\left(\sigma_{t+1}-\frac{\alpha_{t+1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t)$
+$:={\textrm{Inversion}}(x_{t},t\rightarrow t+1).$ (4)
+III-C Multi-view diffusion models
 After being trained on a modest amount of 3D data, diffusion models can generate 3D-consistent multi-view, known as multi-view diffusion models.
@@
 This paper directly utilizes these pre-trained multi-view diffusion models without any further fine-tuning.
-For more details, please refer to the original papers and the source code in our supplementary materials.
-Notably, such multi-view diffusion models cannot provide a sufficient number of consistent multi-view images for 3D reconstruction directly, as shown in Fig.
-7(a).
-Given that the multi-view diffusion in [38, 39, 40, 41] often incorporate with additional 3D priors and do not align with our setting in Sec.
-4, we choose not to employ them in this paper.
-2.4 Optimization-based algorithms for 3D generation
+Notably, such multi-view diffusion models cannot provide sufficient consistent multi-view images for 3D reconstruction directly (see Fig.
+12(a)).
+Given that the multi-view diffusion models in [59, 60, 61, 62, 64, 63] often incorporate additional 3D priors and do not align with our setting in Sec.
+V, we choose not to employ them in this paper.
+III-D Optimization-based algorithms for 3D generation
 Built upon (multi-view) diffusion models, optimization-based methods (see Sec.
-6 for a review) aim to generate 3D content in a zero-shot manner.
+II for a review) aim to generate 3D content in a zero-shot manner.
 Among them, score distillation sampling (SDS) [1, 2] is the most representative and popular approach.
@@
 (1) and updating the 3D parameters $\theta$ by
-$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi):=\mathbb{E}_{t}\left[w(t)(\epsilon_{\phi}(\alpha_{t}x+\sigma_{t}\epsilon,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (5)
-where $w(t)$ is a fixed weighting function and we omit the dependency on the prompt $y$ and camera $c$ for simplicity.
+$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi)$
+::=:= $\mathbb{E}_{t}\left[w(t)(\epsilon_{\phi}(\alpha_{t}x+\sigma_{t}\epsilon,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (5)
+where $w(t)$ is a fixed weighting function, we omit the dependency on the prompt $y$ and camera $c$ for simplicity.
 Notably, by reparameterization [4] according to Eq.
@@
 $\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi)=\mathbb{E}_{t}\left[\frac{w(t)\alpha_{t}}{\sigma_{t}}(x-\hat{x}_{0}^{t})\frac{\partial x}{\partial\theta}\right].$ (6)
-3 Score-based iterative reconstruction
-We present the general framework of score-based iterative reconstruction (SIR), an efficient and versatile algorithm combining both the advantages of differentiable 3D reconstruction and optimization methods in Sec.
-3.1.
-SIR capitalizes on the efficiency of 3D reconstruction and aims to minimize the number of function evaluations (NFEs) typically required by existing methods while ensuring comparable quality through iterations.
-For better reconstruction, we introduce a carefully designed diffusion-based process to produce refined multi-view images as ground truth in Sec.
-3.2.
-3.1 Mimicking 3D reconstruction to reduce the NFEs
-Existing zero-shot 3D generation algorithms [1, 2, 4, 8] iteratively optimize the 3D parameters based on a 2D diffusion (see Eq.
-(5) for SDS as an example).
-Consequently, these methods necessitate high NFEs, presenting a critical challenge to computational efficiency.
-Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of differentiable 3D reconstruction could substantially decrease the overall NFEs required.
+IV Score-based iterative reconstruction
+We first analyze factors contributing to the efficiency bottleneck of existing work in Sec.
+IV-A, motivating score-based iterative reconstruction (SIR), an efficient and versatile algorithm combining the advantages of differentiable 3D reconstruction and optimization methods in Sec.
+IV-B.
+We introduce a carefully designed diffusion-based process to produce refined multi-view images as ground truth for better reconstruction in Sec.
+IV-C.
+Besides, we demonstrate how SIR facilitates the optimization of 3D content directly within pixel space in Sec.
+IV-D.
+IV-A Efficiency bottleneck of existing work
+Figure 4: Time proportion in SDS optimization. We record the time proportions of all components in SDS on 3DGS. The bottleneck lies in the large NFEs and updating in latent space.
+As a motivation of SIR, we analyze the efficiency bottleneck of existing optimization-based methods [1, 2, 4, 9, 17, 18, 7].
+We take the widely adopted SDS [1] as a representative example and our analysis also applies to other algorithms.
+On the one hand, according to Eq.
+(5), SDS iteratively optimizes the 3D parameters based on a 2D diffusion.
+It necessitates high NFEs because it requires a forward pass of the 2D diffusion in each update of the 3D parameters.
+On the other hand, when employed within the latent diffusion model (LDM) [23] upon a variational auto-encoder (VAE) [83, 55], SDS computes the loss in latent space.
+This process requires backpropagation through the VAE encoder, further lowering generation efficiency.
+These arguments are validated by the empirical results in Fig.
+4 quantitatively.
+As for the latent space optimization, we emphasize that although SDS has a corresponding data-prediction form in Eq.
+(6), it is nontrivial to map the predicted $x_{0}^{(t)}$ back to pixel space through the VAE decoder for loss calculation.
+This is because the single-step prediction in Eq.
+(2) yields poor samples in the pixel space, leading to suboptimal 3D results (See experiments in Tab.
+11 of Sec.
+VI).
+IV-B Mimicking 3D reconstruction to reduce the NFEs
+Motivated by the analysis of SDS, we propose SIR to reduce NFEs and enable optimization in pixel space to enhance efficiency.
+Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of differentiable 3D reconstruction could substantially decrease the overall NFEs required and enable optimization in pixel space.
 However, 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [6, 67, 69, 68] (see Sec.
-6 for a review and discussion).
+II for a review and discussion).
 Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple iterations of optimizing dubbed score-based iterative reconstruction (SIR), detailed as follows.
@@
 In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given initial 3D parameters $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec.
-4) and employ rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:
+V) and employ rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:
 $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)}).$ (7)
-For simplicity, let $x^{(k)}$ be a vector by flattening all images and concatenating them together and so do any subsequent set of multi-view images.
+For simplicity, let $x^{(k)}$ be a vector by flattening all images and concatenating them together, and so do any subsequent set of multi-view images.
 As detailed in Sec.
-3.2, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [42, 6, 44, 43].
+IV-C, $x^{(k)}$ is then refined via a pre-trained multi-view diffusion model [67, 6, 69, 68].
 The outcome, denoted as $\hat{x}^{(k)}$, serves as the ground truth for reconstruction in the current iteration.
 In particular, starting from $\theta_{0}^{(k)}$, keeping the camera poses $c^{(k)}$ unchanged, we optimize the 3D parameters w.r.t.
-the following reconstruction loss for $I^{(k)}$ steps:
+the following reconstruction loss for $I$ steps:
 $\mathcal{J}_{\textrm{SIR}}(\theta;c^{(k)},\hat{x}^{(k)})=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|,$ (8)
 […]
-Here we assume the rendered images and those sampled by the diffusion model are the same dimensional for simplicity.
-In Appendix B.1 we will further discuss how to deal with the latent diffusion model (LDM) [17].
+Here we assume the rendered images and those sampled by the diffusion model have the same dimensional for simplicity.
+In Sec.
+IV-D we will further discuss how to deal with the latent diffusion model (LDM) [23].
 Before that, we discuss how to obtain 3D-consistent and high quality $\hat{x}^{(k)}$ from $x^{(k)}$ for reconstruction in Sec.
-3.2.
-3.2 Refining multi-view images for reconstruction by diffusion
-We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process (see Fig. 3b).
-Figure 3: Overview of DDIM reconstruction and our refinement.
-(a) The inversion process can preserve the 3D consistency of input images, and the input images can be recovered through the corresponding DDIM sampler.
-(b) After the carefully designed refinement of $x$ by Eq.
-(9) and Eq.
-(10), the resulting $\hat{x}$ possesses enough high quality for 3D reconstruction.
-The diffusion model has two equivalent forward processes [16]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq.
+IV-C.
+Figure 5: The hybrid forward process is more efficient than DDIM inversion and generates better samples than noise-adding.
+We present the final results and sampling time on 20 iterations of SIR for three forward processes.
+Notably, the Noise-adding process may generate artifacts that contain unexpected elements compared to the input.
+IV-C Refining multi-view images for reconstruction by diffusion
+Algorithm 1 Score-based iterative reconstruction (SIR)
+1:  Input: The number of iterations $K$, an initial 3D object $\theta_{0}^{(0)}$, the number of reconstruction steps $I$ and a set of camera poses $\{c^{(k)}\}_{k=0}^{K-1}$.
+2:  Output: A final 3D content $\theta^{(K)}_{0}$.
+3:  for $k$ from 00 to $K-1$ do
+4:     Render $N$ images $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)})$
+5:     Obtain noisy $\tilde{x}^{(k)}$ from forward process for $x^{(k)}$
+6:     Obtain refined $\hat{x}^{(k)}$ from sampling process for $\tilde{x}^{(k)}$
+7:     for $i$ from 00 to $I-1$ do
+8:        Compute loss $L=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|$
+9:        Compute the gradient $\nabla L$, update $\theta^{(k)}_{i}$ to $\theta^{(k)}_{i+1}$
+10:     end for
+11:  end for
+We treat $x^{(k)}$ as noise-free (i.e., at time 00 of the diffusion process) but low-quality images, refined by a forward process followed by a sampling process.
+The diffusion model has two theoretically equivalent forward processes [22]: the noise-adding process modeled by stochastic differential equations (SDEs) in Eq.
 (1) and the inversion process based on the probability flow ordinary differential equations (ODEs) in Eq.
 (4).
-In comparison, the noise-adding process is more efficient without function evaluation, and it aids in eliminating low-quality details through high-frequency noise.
-Conversely, the inversion process can preserve the 3D consistency and overall information of the current 3D object (see Fig.
-3a).
-In addition, observations from existing experiments  [55] indicate that diffusion models tend to generate details closer to time 00 and overall structure as they approach time $T$.
-Inspired by this, we design a hybrid forward process that initially adds noise to time $t_{1}^{(k)}$ and then perform DDIM inversion to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise..
+In comparison, the noise-adding process is more efficient without function evaluation, but sampling after the noise-adding process may produce unexpected artifacts (see Fig.
+5).
+The inversion process can better preserve the 3D consistency and overall information of the current 3D object, but it necessitates more NFEs with low efficiency.
+We carefully design a hybrid forward process that initially adds noise and then performs DDIM inversion.
+Compared to common noise-adding and inversion processes, the hybrid forward process achieves a better balance in quality and efficiency, as Fig.
+5 shows.
+Specifically, the hybrid forward process adds noise to time $t_{1}^{(k)}$ first and then performs DDIM inversion [80] to time $t_{2}^{(k)}$, where $t_{1}^{(k)}\in(0,T)$ and $t_{2}^{(k)}\in[t_{1}^{(k)},T)$ are hyperparameters [^2] Note that if $t^{(k)}_{1}=t^{(k)}_{2}$, the process is purely adding noise..
 In contrast to random sampled $t_{2}^{(k)}$ in existing algorithms including SDS [1], we adopt a linearly decreased schedule for $t_{2}^{(k)}$ as $k$ progresses, detailed in Sec.
-4.
+V.
 Formally, the process is defined as:
@@
 […]
-Note that generally $x^{(k)}$ does not follow the model distribution defined by the diffusion and the resulting $\tilde{x}^{(k)}$ do not strictly adhere to the marginal distribution at the corresponding time.
+Note that generally $x^{(k)}$ does not follow the model distribution defined by the diffusion and the resulting $\tilde{x}^{(k)}$ does not strictly adhere to the marginal distribution at the corresponding time.
 Nevertheless, we still use existing sampling algorithms to obtain refined images $\hat{x}^{(k)}$ from $\tilde{x}^{(k)}$ as follows:
@@
 The optimal value of $\eta$ depends on the base model and we search for the best one in $\{0,0.5,1\}$ as detailed in Tab.
-1 in Appendix A.
+I.
 The whole process of SIR is presented in Algorithm 1.
-Compared with existing generation methods such as SDS, SIR benefits from reconstruction, i.e.
-leverages several NFEs for multiple updates of 3D parameters, and lowers the NFEs in total (see Fig.
-4).
-1
-2Input: The number of iterations $K$, an initial 3D object $\theta_{0}^{(0)}$, a set of reconstruction steps $\{I^{(k)}\}_{k=0}^{K-1}$ and a set of camera poses $\{c^{(k)}\}_{k=0}^{K-1}$.
-3 Output: A final 3D content $\theta^{(K)}_{0}$. for $k$ from 00 to $K-1$ do
-4       Render $N$ images $x^{(k)}=g(\theta^{(k)}_{0},c^{(k)})$ ;
-5       Obtain noisy $\tilde{x}^{(k)}$ from forward process for $x^{(k)}$ using Eq. (9);
-6       Obtain refined $\hat{x}^{(k)}$ from sampling process for $\tilde{x}^{(k)}$ using Eq. (10);
-7       for $i$ from 00 to $I^{(k)}-1$ do
-8             Compute loss $L=\|g(\theta,c^{(k)})-\hat{x}^{(k)}\|$ using Eq. (8);
-9             Compute the gradient $\nabla L$, update $\theta^{(k)}_{i}$ to $\theta^{(k)}_{i+1}$;
-10
-11       end for
-12      Obtain $\theta^{(k+1)}_{0}=\theta^{(k)}_{I^{(k)}}$.
-13 end for
-Algorithm 1 Score-based iterative reconstruction
-4 MicroDreamer
-In addition to our algorithm SIR, we carefully analyze and design other critical elements in the 3D training and diffusion. Our comprehensive system is called MicroDreamer to highlight its efficiency.
+Compared with existing methods [1, 7, 9, 15] dedicated to enhancing 3D generation quality, SIR benefits from reconstruction and lowers the NFEs in total, consequently improving generation efficiency.
+(see Fig.
+7 for the comparison to SDS).
+We visualize the optimization process of SIR in Fig.
+6.
+With the increase in the number of iterations, the visual quality of the 3D content improves, validating the effectiveness of our iterative reconstruction.
+Figure 6: Visualization of the optimization process in SIR. The visual quality of the 3D samples increases along with the iterations.
+IV-D Enabling optimization in pixel space
+Within the widely adopted framework of LDM [23], the rendered images are mapped through an encoder $\mathcal{E}$ [54, 55] to a latent space, where the loss function is calculated.
+Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation.
+In such a case, the SIR loss in latent space is given by:
+$\mathcal{J}_{\textrm{SIR-latent}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|\mathcal{E}(g(\theta,c^{(k)}))-\hat{x}^{(k)}\|.$ (11)
+An alternative approach maps the diffusion output back to pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space.
+This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
+As discussed in Sec.
+IV-A, previous optimization-based methods like SDS fail to be effective when applied in pixel space.
+In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space.
+The SIR loss in pixel space is formalized as:
+$\mathcal{J}_{\textrm{SIR-pixel}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|g(\theta,c^{(k)})-\mathcal{D}(\hat{x}^{(k)})\|,$ (12)
+which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as the analysis results shown in Fig. 10. It is set as the default loss throughout the paper.
+V MicroDreamer
+We provide a general configuration in the 3D training and diffusion for the SIR algorithm. The comprehensive system is called MicroDreamer to highlight its efficiency.
 3D initialization and camera views.
@@
 (8) by several steps (see Tab.
-1 in Appendix A for detailed values) to update the 3D parameter, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion model.
-The camera views are uniformly sampled following the corresponding baselines [9, 56] except that the azimuth angles of different views in the same iteration are evenly distributed.
-In this setting, we are unable to directly utilize the diffusion model in [38, 39, 40] due to their fixed camera views conditioned generation.
+I for detailed values) to update the 3D parameters, where $\hat{x}$ represents the images sampled from random noise via the pre-trained multi-view diffusion models.
+The camera views are uniformly sampled following the corresponding baselines [8, 84] except the azimuth angles of different views in the same iteration are evenly distributed.
+In this setting, we are unable to directly utilize the diffusion model in [59, 60, 62, 63] due to their fixed camera views conditioned generation.
 Annealed time schedule.
@@
 Our preliminary experiments in Fig.
-7(c) suggest that a linearly annealed schedule is sufficient.
+12(c) suggest that a linearly annealed schedule is sufficient.
 The endpoints of the schedule depend on the diffusion model, detailed in Tab.
-1 in Appendix A.
-See Appendix C for the schedule of $t_{1}$.
-Reference image loss for image-to-3D.
-A reference image is available in image-to-3D, which is regarded as the ground truth front view of the 3D object on Stable Zero123 [6, 43] following DreamGaussian [9].
-In this way, we add the reference loss in the same form as the reconstruction loss with Eq.
-(8) in each iteration of training.
-The weight set for the reference loss can be seen in Tab.
-1.
-3DGS settings and mesh refinement.
-For simplicity, our approach on 3DGS largely adheres to the settings in DreamGaussian [9] unless specified.
-We incorporate a densify and prune procedure at every 50 updates during the initial 200 updates.
-We follow the mesh extraction method in DreamGaussian  [9] and employ a threshold value of $0.2$ in the marching cube algorithm [57, 9].
-We utilize SIR to optimize the exported mesh texture in 3 iterations, with 15 steps per iteration, and we use a noise-adding process and DDIM with $\eta=0$ for simplicity.
-5 Experiment
-We present the results on NeRF, 3DGS, and ablation sequentially.
-For a fair comparison, we utilize the widely used CLIP similarity [45] for quantitative comparison following DreamGaussian [9].
-We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0.
-For all methods, the corresponding images are rendered from NeRF in Sec.
-5.1 and mesh in Sec.
-5.2.
-We report the average CLIP similarities between these images and the reference image or text.
-See Appendix D for details of the checkpoint and codebase.
-5.1 Results on NeRF
-We apply SIR algorithm on NeRF [35, 36, 56] leveraging three multi-view diffusion model from MVDream [42], Stable Zero123 [43, 6], and ImageDream [44].
-For each model, we selected 6 input prompts from a widely used codebase [56] for testing, calculated NFEs, and recorded the change of the average CLIP similarity during the generation process.
-Compared with SDS, SIR lowers the total NFEs while generating a 3D object and can accelerate the generation process 5-20 times, as shown in Fig.
-4.
-Besides, SIR can generate NeRF of high visual quality, as detailed shown in Appendix E, and the conclusion agrees with the results in Fig.
-4.
-(a) Comparison on MVDream
-(b) Comparison on Stable Zero123
-(c) Comparison on ImageDream
-Figure 4: Comparison of SIR and SDS on NeRF. The change of the CLIP similarity in the generation process and final NFEs on different models. SIR achieves a 5-20 times acceleration and lowers NFEs.
-5.2 Results on 3D Gaussian splatting
-Qualitative comparisons.
-We show the generated 3D meshes between the fastest optimization-based baseline DreamGaussian [9] and ours in Fig.
-5.
-MicroDreamer halves the generation time of DreamGaussian while retaining competitive performance.
-We provide more visualization results of our methods in Appendix E.
-Figure 5: Qualitative comparisons on mesh from 3DGS. MicroDreamer retains a comparable performance in terms of the visual quality of the mesh with higher efficiency.
-Quantitative comparisons.
-We compare MicroDreamer with seven competitive baselines including Point-E [24], Shap-E [23], One-2-3-45 [31], TriplanceGaussian [33], Wonder3D [39], LGM [30] and DreamGaussian [9].
-We record the optimization time  [^3] The detailed time recording method is provided in Appendix D.
-for each model on a single NVIDIA A100 (80GB) GPU, and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works [31, 38, 9], as shown in Fig.
-7.
-As a zero-shot method, MicroDreamer is on par with speed compared to feed-forward methods, faster than DreamGaussian, with a very competitive CLIP similarity.
-We provide the time record and detailed tables in Appendix F.
-Notably, the sample quality of MicroDreamer can be improved as the multi-view diffusion evolves.
-See detailed limitations and future work in Sec.
-7 and Appendix E.
-Figure 6: Quantitative comparisons. Our MicroDreamer achieves an excellent efficiency-effectiveness balance in terms of CLIP similarity. The zero-shot methods are highlighted in red.
-[…]
-(a) Iterative reconstruction.
-(b) Camera views.
-(c) Time schedule.
-(d) Forward process.
-Figure 7: Ablation study. (a) Iterative reconstruction is necessary. (b) More camera views benefit. (c) A linear schedule is sufficient. (d) A hybrid forward process can be effective.
-5.3 Ablation study
-Ablation on iterative reconstruction.
-Fig.
-7(a) shows the necessity of iterative reconstruction.
-The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise.
-Although increasing $K$ improves quality, we set $K\leq 30$ for efficiency.
-Ablation on number of camera poses.
-SIR necessitates more than 1 camera pose for reconstruction, as shown in Fig.
-7(b).
-The performance increases with more camera poses.
-We choose $6$ on 3DGS with Stable Zero123 [43] model to balance efficiency and quality.
-See Tab.
-1 for values in other settings.
-Ablation on time schedule.
-We compare our linear schedule, the random schedule employed in SDS [1], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.8T-0.2T)+0.2T$.
-As shown in Fig.
-7(c), the schedules perform similarly.
-For simplicity and efficiency, we use the linear schedule by default.
-Ablation on forward process.
-Fig.
-7(d) compares three forward processes on 3DGS with the Stable Zero123 [43] model.
-The hybrid strategy performs better than noise-adding and is more efficient than inversion as discussed in Sec.
-3.2.
-See Tab.
-1 in Appendix A for other settings.
-6 Related work
-Optimization-based 3D generation.
-Built upon text-to-image diffusion models, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 58, 59, 60] usually generate 3D objects without additional 3D data.
-Among them, the most relevant line of work [1, 8, 10, 4, 2] proposes various distillation algorithms.
-Besides SDS [1], ProlificDreamer [8] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference.
-LucidDreamer [10] incorporates DDIM inversion to strengthen the forward process.
-Though promising, such methods suffer from inefficiency problems due to large NFEs.
-Distinct from existing algorithms, SIR mimics 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion.
-Our experiments demonstrate that SIR is 5-20 times faster than SDS [1] on NeRF and twice as fast as the most competitive baseline DreamGaussian [9] on meshes.
-Feed-forward 3D generation.
-Contrary to optimization-based methods, feed-forward methods [20, 21, 22, 23, 24, 25, 26, 27, 28, 61, 29, 31, 32, 33, 34, 62, 63, 64, 65] elect to use large-scale 3D datasets [66, 67] training to achieve a process that can directly generate 3D objects.
-Early methods are characterized by their speed but often produce lower-quality 3D structures with simple textures.
-Several recent methods, utilizing more data, have exhibited possibilities.
-Large reconstruction model (LRM) [68, 26] demonstrates the feasibility of training a Transformer to achieve a NeRF from a single image.
-Instant3D [69] further expands on LRM enabling it to accept multi-view input and reconstruct it into a NeRF.
-Remarkably, MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [45, 9].
-Multi-view prediction based 3D generation.
-There is also a line of work [38, 39, 40, 41] dedicated to enhancing the output of multi-view diffusion models by training on 3D datasets to reconstruct 3D objects using a single reconstruction process without iteration.
-These methods like Wonder3D[39] typically rely on 3D prior knowledge and are limited to specific 3D representations with a long reconstruction time.
-In contrast, thanks to the iterative process in SIR, MicroDreamer applies to various 3D representations and is more efficient (see Fig.
-7 for a direct comparison).
-7 Conclusion
-We introduce SIR, an efficient and general algorithm combining both the strengths of 3D reconstruction and iterative optimization to reduce total NFEs in zero-shot 3D generation.
-SIR achieves a 5 to 20 times speed increase for NeRF generation compared to SDS.
-Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest zero-shot methods and matching the speed of feed-forward approaches with a competitive generation quality.
-Limitations and broader impact.
-There is potential for further improving MicroDreamer’s efficiency via employing consistency models [70, 71] or alternative sampling models that require fewer steps [72, 73].
-Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion.
-Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
-Also, MicroDreamer can provide convenience for artists to create 3D assets.
-However, as a generative approach, our method may also be used for fabricating data and news, which warrants careful consideration.
-8 Acknowledgement
-Thanks Cheng Lu for the discussion when he was a PhD student at Tsinghua University.
-Appendix A Key hyperparameters of MicroDreamer
-Table 1: Key hyperparameters of MicroDreamer on three base diffusion models.
+I.
+TABLE I: Key hyperparameters of MicroDreamer on three base diffusion models.
 All models are employed to generate NeRF and the last two are employed to generate 3DGS and mesh.
-The hyperparameters are shared across the two settings by default.
+By default, the hyperparameters are shared across the two 3D representations.
 Otherwise, those for 3DGS are shown in brackets.
-$T$ is end time of diffusion.
-Model Select MVDream [42] Stable Zero123 [6, 43] ImageDream [44]
+$T$ is the end time of diffusion.
+Model Select MVDream [67] Stable Zero123 [68] ImageDream [69]
 Diffusion
 CFG [85] 7.5 3.0 3.0 (2.0)
-Forward process hybrid hybrid hybrid (random noise only)
-Time schedule of $t_{2}$ $0.8T\rightarrow 0.5T$ $0.8T\rightarrow 0.2T$ $0.8T\rightarrow 0.6T$ ($0.8T\rightarrow 0.4T$)
+Forward process hybrid hybrid hybrid
+Time schedule of $t_{2}$ $0.8T\rightarrow 0.5T$ $0.8T\rightarrow 0.2T$ ($0.9T\rightarrow 0.2T$) $0.8T\rightarrow 0.6T$ ($0.8T\rightarrow 0.4T$)
+Time schedule of $t_{1}$ $t_{1}=t^{2}_{2}/T$ $t_{1}=0.6t_{2}$ $t_{1}=t^{2}_{2}/T$ ($t_{1}=0.6t_{2}$)
 Sampling process DDIM, $\eta=0.0$ DDIM, $\eta=0.5$ DDIM, $\eta=1.0$
@@
 # iterations $K$ 50 30 (20 or 30) 30
-# reconstuction steps $I^{(k)}$ 15 15 15
+# reconstruction steps $I$ 15 15 15
 Loss type $\ell_{1}$ $\ell_{1}$ $\ell_{1}$
-Reference color loss - 0.1 (0.3) 0
-Reference opacity loss - 0.001 (0.01) 0
-Appendix B The impact of latent diffusion model on 3D generation
-Instead of applying the diffusion model directly in the pixel space to generate images, existing research [17] tends to use a VAE [75, 76] encoder to compress high-dimensional image data into a low-dimensional latent space first.
-Then a diffusion process is employed within this latent space to generate new images, which are finally decoded back by a VAE decoder to the pixel space to achieve high-quality images.
-We refer to this type of diffusion model as the latent diffusion model (LDM).
-The framework of LDM is adopted widely [6, 42, 44, 43], which leads to a significant impact on 3D generation.
-B.1 SIR enabling optimization in the pixel space
-Within the framework of LDM [17], the rendered images are mapped through an encoder $\mathcal{E}$ [75, 76] to a latent space, where the loss function is calculated.
-Consequently, the gradients must be propagated back through the encoder, further reducing the efficiency of 3D object generation.
-In such a case, the SIR loss in the latent space is given by:
-$\mathcal{J}_{\textrm{SIR-latent}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|\mathcal{E}(g(\theta,c^{(k)}))-\hat{x}^{(k)}\|.$ (11)
-An alternative approach maps the diffusion output back to the pixel space via the corresponding decoder $\mathcal{D}$, allowing for a similar loss function to be defined directly in pixel space.
-This method enables direct updates to 3D parameters without passing gradients through the encoder, thus enhancing efficiency.
-However, we argue that this is not feasible for existing optimization methods [1, 2, 8] such as SDS.
-Even reparameterized as the data predicting form in Eq.
-(6), the pixel-level quality of the images obtained by SDS is poor for 3D optimization (see Appendix B.2).
-In contrast, SIR achieves higher quality generation results through a carefully designed refinement process, thereby enabling optimization in pixel space.
-The SIR loss in the pixel space is formalized as:
-$\mathcal{J}_{\textrm{SIR-pixel}}(\theta;c^{(k)},\hat{x}^{(k)})$ $=\|g(\theta,c^{(k)})-\mathcal{D}(\hat{x}^{(k)})\|,$ (12)
-which is about 2-3 times faster than Eq. (11) for 3D parameter optimization, as the ablation results shown in Fig. 8. It is set as the default loss throughout the paper.
+Ref. color loss - 0.1 (0.3) 0
+Ref. opacity loss - 0.001 (0.01) 0
+Reference image loss for image-to-3D.
+A reference image is available in image-to-3D, which is regarded as the ground truth front view of the 3D object on Stable Zero123 [68] following DreamGaussian [8].
+In this way, we add the reference loss in the same form as the reconstruction loss with Eq.
+(8) in each training iteration.
+The weight set for the reference loss can be seen in Tab.
+I.
+3DGS settings and mesh refinement.
+For simplicity, our approach on 3DGS largely adheres to the settings in DreamGaussian [8] unless specified.
+We incorporate a densify and prune procedure at every 100 updates during the initial 300 updates.
+At the end of optimization, we do a last prune and remove potential white Gaussians with scales larger than 0.01.
+We follow the mesh extraction method in LGM [43] and employ a threshold value of $5$ in the marching cube algorithm [86].
+We utilize SIR to optimize the exported mesh texture for one iteration with 30 steps, and we use a noise-adding process and DDIM with $\eta=0$ for simplicity.
+VI Experiments
+We present the experimental details and results on NeRF, 3DGS, and ablation sequentially.
+VI-A Experimental details
+We present key hyperparameters of MicroDreamer in Tab.
+I.
+For implementing SIR and SDS on NeRF, we choose the popularly used framework threestudio [84].
+For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [8].
+Most hyperparameters are not sensitive to the results, and we follow the previous framework [84, 8] for these settings.
+For important hyperparameters, including the number of iterations, camera views, and the time schedule, see Sec.
+VI-D for ablation.
+See Tab.
+II for details about the URL of the codebases and checkpoints we used in this paper.
+For a fair comparison, we utilize the widely used CLIP similarity [71] for quantitative comparison following DreamGaussian [8].
+We consider 8 views at 0 elevation and evenly distributed in azimuth angles starting from 0.
+For all methods, the corresponding images are rendered from NeRF in Sec.
+VI-B and mesh in Sec.
+VI-C.
+We report the average CLIP similarities between these images and the reference image or text.
+TABLE II: Codebases and checkpoints. We provide URLs for the open-source assets we used in this paper.
+Model URL
+Codebases
+[84] https://github.com/threestudio-project/threestudio
+[8] https://github.com/dreamgaussian/dreamgaussian
+Checkpoints
+[87] https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k
+[67] https://github.com/bytedance/MVDream-threestudio
+[68] https://huggingface.co/stabilityai/stable-zero123
+[69] https://github.com/bytedance/ImageDream
+Baselines
+[30] https://github.com/openai/point-e
+[29] https://github.com/openai/shap-e
+[39] https://github.com/One-2-3-45/One-2-3-45
+[37] https://github.com/VAST-AI-Research/TriplaneGaussian
+[60] https://github.com/xxlong0/Wonder3D
+[43] https://github.com/3DTopia/LGM
+[70] https://github.com/3DTopia/OpenLRM
+(a) Comparion on MVDream [67]
+(b) Comparion on Stable Zero123 [68]
+(c) Comparion on ImageDream [69]
+Figure 7: Comparison of SIR and SDS on NeRF.
+We plot the curve of the CLIP similarity in the generation process and final NFEs on different models.
+SIR lowers the NFEs and achieves a 5-20 times acceleration to achieve a competitive quality.
+Figure 8: Qualitative comparison on NeRF. SIR can generate NeRF of higher visual quality than SDS in a short time.
+VI-B Results on NeRF
+We apply SIR algorithm on NeRF [57, 58, 84] leveraging three multi-view diffusion models from MVDream [67], Stable Zero123 [68], and ImageDream [69].
+For each model, we selected 6 input prompts from a widely used codebase [84] for testing, calculated NFEs, and recorded the average CLIP similarity during the generation process.
+Compared with SDS, SIR lowers the total NFEs and accelerates the generation process 5-20 times while holding a competitive CLIP similarity, as shown in Fig.
+7.
+To provide a comprehensive comparison, we qualitatively analyze the generation results of SDS and SIR under identical time conditions, as illustrated in Fig.
+8.
+Notably, at the moment of convergence in SIR, we find that SDS has not achieved satisfactory results, such as producing ambiguous outcomes.
+These observations support the quantitative results, demonstrating that SIR is general and more efficient.
+VI-C Results on 3D Gaussian splatting
+Qualitative comparisons.
+In Fig.
+9, we present the generated 3D meshes comparing the fastest optimization-based baseline, DreamGaussian [8], with our method.
+MicroDreamer reduces DreamGaussian’s generation time by half while exhibiting superior performance, including improved texture and enhanced geometric structure.
+Figure 9: Qualitative comparisons on mesh from 3DGS.
+MicroDreamer produces superior meshes, characterized by enhanced texture and reduced geometric artifacts, with greater efficiency than DreamGaussian.
+Quantitative comparisons.
+We compare MicroDreamer with eight competitive baselines including Point-E [30], Shap-E [29], One-2-3-45 [39], TriplaneGaussian [^3] As TriplaneGaussian has no official mesh export code, we apply the mesh exported code from LGM [43] for it.
+[37], Wonder3D [60], LGM [43], Open-LRM [70] and DreamGaussian [8].
+We record the generation time for each model on a single NVIDIA A100 (80GB) GPU and compute the average CLIP similarity for mesh on a test dataset consisting of 87 images collected from previous works [39, 59, 8], as shown in Tab.
+III.
+MicroDreamer generates significantly better 3D content than DreamGaussian, as indicated by the standard deviation across multiple runs, and has higher efficiency.
+In addition, MicroDreamer is on par with speed compared to feed-forward methods [70] trained on a substantial amount of 3D data and has a very competitive CLIP similarity.
+In conclusion, all those results suggest that SIR is a promising approach for efficient 3D generation.
+TABLE III: Quantitative comparisons.
+MicroDreamer significantly outperforms the strong optimization-based baseline DreamGaussian in quality and efficiency and remains competitive with feed-forward methods.
+All results are averaged over three runs.
+Method CLIP similarity $\uparrow$ Generation time $\downarrow$
+Point-E [30] $0.566\pm 0.0011$ $\sim 24$s
+Shap-E [29] $0.626\pm 0.0030$ $\sim 5$s
+One-2-3-45 [39] $0.617\pm 0.0025$ $\sim 42$s
+Wonder3D [60] $0.696\pm 0.0017$ $\sim 170$s
+TriplaneGaussian [37] $0.691\pm 0.0016$ $\sim 7$s
+LGM [43] $0.700\pm 0.0017$ $\sim 3$s
+Open-LRM-mix-base-1.1[70] $0.704\pm 0.0000$ $\sim 23$s
+DreamGaussian [8] $0.692\pm 0.0015$ $\sim 30+10$s
+DreamGaussian-300 iter [8] $0.641\pm 0.0032$ $\sim 18+10$s
+MicroDreamer-20 iter (Ours) $0.697\pm 0.0009$ $\sim 18+3$s
+MicroDreamer-30 iter (Ours) $0.711\pm 0.0007$ $\sim 26+3$s
+VI-D Analysis and ablation study
+(a) On Stable Zero123[68].
+(b) On ImageDream [69].
 Figure 10: Optimization in pixel space accelerates generation.
-B.2 Faliure of applying SDS in pixel space.
-We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq.
+(a) Results of single-step prediction using Eq. (2) for SDS.
+(b) Results of applying SDS in pixel space.
+Figure 11: Failure of applying SDS in pixel space. SDS optimizing in pixel space generates poor results.
+(a) Iterative reconstruction.
+(b) Camera views.
+(c) Time schedule.
+(d) Forward process.
+Figure 12: Detailed analyses. (a) Iterative reconstruction is necessary. (b) More camera views benefit. (c) A linear schedule is sufficient. (d) A hybrid forward process can be effective.
+Analysis on optimization space.
+We implement SIR-latent on NeRF and compare its time consumption with SIR-pixel, as illustrated in Fig.
+12a-b.
+We can see that SIR-pixel achieves 2-3 times more acceleration than SIR-latent, verifying the benefits of SIR in enabling optimization in pixel space.
+Contrary to SIR, we argue that SDS optimizing in pixel space is not effective even if it has the data predicting form by reparameterization as presented in Eq.
 (6).
+If we attempt to employ SDS for optimization in pixel space, i.e.
+optimizing 3D objects with 2D images like those generated by Eq.
+(2) in Fig.
+11a, the resulting 3D content would be of subpar quality, as the results shown in Fig.
+11b.
+Ablation on iterative reconstruction.
 Fig.
-9 shows the poor quality for single-step prediction from the diffusion model using Eq.
-(2).
-Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e.
-optimizing 3D objects with 2D images like those generated in Fig.
-9, the resulting 3D content would be of subpar quality, as the results presented in Fig.
-10.
-Figure 9: Results of single-step prediction.
-For each pair of images, the left is the input image and the right is the output of single-step prediction.
-The poor quality of the output makes it unsuitable for 3D optimization.
-Figure 10: Results of applying SDS in pixel space. SDS optimizing in pixel space can not be effective.
-Appendix C Time schedule of $t_{1}$
-The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models.
-In our earlier exploration, we adopted $t_{1}=t^{2}_{2}/T$ for SIR based on MVDream [42] and ImageDream [44] on NeRF, and $t_{1}=0.6t_{2}$ for SIR based on Stable Zero123 [43, 6] on both NeRF and 3DGS.
-Notably, SIR based on ImageDream on 3DGS opted for $t_{1}=t_{2}$, i.e., adding noise for the forward process.
-As an advanced multi-view diffusion model, ImageDream can generate images with enhanced 3D consistency, obviating the need for inversion to maintain consistency.
+12(a) shows the necessity of iterative reconstruction.
+The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise.
+When $K>30$, the generation quality doesn’t improve significantly.
+For efficiency, we set $K=20$ or $30$.
+Ablation on number of camera poses.
+SIR necessitates more than 1 camera pose for reconstruction, as shown in Fig.
+12(b), and the performance increases with more camera poses.
+We choose $6$ on 3DGS with Stable Zero123 [68] model to balance efficiency and quality.
+See Tab.
+I for values in other settings.
+Ablation on time schedule.
+We compare our linear schedule, the random schedule employed in SDS [1], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.9T-0.2T)+0.2T$.
+As shown in Fig.
+12(c), the schedules perform similarly.
+For simplicity and efficiency, we use the linear schedule by default.
+Ablation on forward process.
 Fig.
-11 shows some results of SIR on ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
-Figure 11: Results of SIR based on ImageDream on 3DGS. These results images are rendered from 3DGS with no refinement stage.
-Appendix D More implementation details
-Codebase.
-For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [56].
-For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [9].
-The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [45, 77, 78].
-In Sec.
-5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub.
-See Tab.
-2 for details about the URL and licenses of the assets we used in this paper.
-Table 2: Codebase. URL and licenses are provided for the open-source assets we used in this paper.
-Details in metric.
-All the results presented in Sec.
-5.2 were on the mesh representation.
-As TriplaneGaussian [33] has no official mesh export code, we apply the mesh exported code from LGM [30] for it.
-For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [9], the time of generating 3D from multi-view images for Wonder3D [39], and the time for meshing through a NeRF for LGM [30], respectively.
-Data used in ablation study.
-We chose seven images collected from the test dataset to perform our ablation study in Fig. 7 and Fig. 8. We will include these images in the submitted supplementary material.
-Safeguards.
-Our method, which requires no additional models and doesn’t release any new pre-trained model, can ensure security due to the inherent safe configuration of the employed 2D multi-view diffusion models [42, 43, 44].
-Appendix E More results
-Results on 3DGS.
-We provide more results on mesh exported from 3DGS for MicroDreamer-20 iters in Fig. 12.
-Figure 12: More results for MicroDreamer-20 iters.
-Results on NeRF.
-In Fig.
-13, we present a comparison of the generated quality between SIR and SDS on NeRF in the same short time under different base diffusion models [42, 43, 6, 44].
-SIR can generate NeRF of high visual quality in a shorter time than SDS, as the results represented.
-Notably, SIR accelerates the generation process in ImageDream more, as it’s an advanced multi-view diffusion model that can produce higher-quality multi-view images and support fewer discretization steps for the diffusion process, lowering NFEs.
-Figure 13: Comparison of the generated quality between SIR and SDS in the same short time on NeRF. SIR can generate NeRF of high visual quality in less time than SDS.
-Limitations.
-In Fig.
-14, we show some of the less successful cases generated by our method.
-Due to the limited ability of the diffusion model, inputs with complex geometric structures such as central hollowing may not result in satisfactory mesh outcomes.
-Furthermore, existing diffusion models struggle to generate the rear view of an object, which often leads to poorly textured back surfaces of the generated meshes.
-These problems may be solved as the quality of the images the multi-view diffusion model generated improves.
-Figure 14: Limitations. Visualization of some less satisfactory cases generated by our methods.
-Appendix F Quantative comparison on 3DGS
-We provide the detailed quantitative comparison table in Tab.
-3.
-For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iters.
-As shown in Tab.
-3, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
-Table 3: Quantitative comparisons. In terms of CLIP similarity, MicroDreamer achieves an excellent balance between efficiency and effectiveness.
-Method CLIP-Similarity $\uparrow$ Generation time $\downarrow$ Extra refine $\downarrow$
-Point-E 0.510 $\sim 24$s -
-Shap-E 0.616 $\sim 5$s -
-One-2-3-45 0.617 $\sim 42$s -
-Wonder3D 0.699 $\sim 30$s $\sim 140$s
-TriplaneGaussian 0.647 $\sim 7$s -
-LGM 0.716 $\sim 3$s $\sim 50$s
-DreamGaussian 0.698 $\sim 30$s $\sim 10$s
-DreamGaussian-300 iters 0.656 $\sim 18$s $\sim 10$s
-MicroDreamer-20 iters (Ours) 0.678 $\sim 18$s $\sim 4$s
-MicroDreamer-30 iters (Ours) 0.687 $\sim 26$s $\sim 4$s
+12(d) compares three different forward processes on 3DGS with the Stable Zero123 [68] model.
+The hybrid strategy performs better than noise-adding.
+As discussed in Sec.
+IV-C, we adopt the hybrid process rather than inversion for efficiency.
+VI-E Limitation
+We show some less successful cases generated by our method in Fig.
+13.
+MicroDreamer faces challenges in generating complex geometric structures such as central hollows and may produce meshes with poor back surface textures.
+These problems may be mitigated as the quality of the generated multi-view images improves.
+Figure 13: Limitation. Visualization of some less satisfactory cases generated by MicroDreamer.
+VII Conclusion
+We introduce SIR, an efficient and general algorithm combining the strengths of 3D reconstruction and iterative optimization to reduce total NFEs and enable optimization in pixel space in optimization-based 3D generation.
+SIR achieves a 5 to 20 times speed increase in NeRF generation compared to SDS.
+Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest optimization-based baseline DreamGaussian in quality and efficiency, and matching the speed of some feed-forward approaches with a competitive generation quality.
+There is potential for further improving MicroDreamer’s efficiency via employing consistency models [88, 89] or alternative sampling models that require fewer steps [90, 91].
+Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion.
+Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
+Furthermore, MicroDreamer can provide artists with the convenience of creating 3D assets.
+However, as a generative approach, our method may also be used for fabricating data and news.
+Moreover, the 3D content generated from input images may infringe on the privacy or copyright of others.
+While automatic detection might mitigate these issues, they warrant careful consideration.